### v1

# Rigidity of the escaping set of polynomial automorphisms of $\mathbb{C}^{2}$

Abstract.
Let $H$ be a polynomial automorphism of $\mathbb{C}^{2}$ of positive entropy and degree $d\geq 2$. We prove that the escaping set $U^{+}$ (or equivalently, the non-escaping set $K^{+}$), of $H$ is rigid under the action of holomorphic automorphisms of $\mathbb{C}^{2}$. Specifically, every holomorphic automorphism of $\mathbb{C}^{2}$ that preserves $U^{+}$ takes the form $L\circ H^{s}$ where $s\in\mathbb{Z}$ and $L$ belongs to a finite cyclic group of affine maps that preserve the escaping set.
Second, note that the sub–level sets $\{G^{+}<c\}$, $c>0$, of the Green’s function $G^{+}$ associated with the map $H$ are canonical examples of Short $\mathbb{C}^{2}$’s. As a consequence of the above theorem, we show that the holomorphic automorphisms of these Short $\mathbb{C}^{2}$’s are affine automorphisms of $\mathbb{C}^{2}$ preserving the escaping set $U^{+}$. Hence, the automorphism group of these Short $\mathbb{C}^{2}$’s are the same for every $c>0$ and is a finite cyclic group.
Key words and phrases:
Hénon map, Escaping set, Automorphism group, Short $\mathbb{C}^{2}$
2020 Mathematics Subject Classification:
Primary: 37F80, 32M18; Secondary: 32Q02, 32T05
1. Introduction
The goal of this article is to study a rigidity property of the escaping set (or equivalently, the non-escaping set) of polynomial automorphisms of $\mathbb{C}^{2}$ of positive topological entropy. By a classical theorem of Friedland–Milnor [14], they are conjugate to a polynomial automorphism $H$ of the form
(1.1) $H(x,y)=H_{m}\circ\cdots\circ H_{1}(x,y)$
where $H_{i}(x,y)=(y,p_{i}(y)-a_{i}x)$, $p_{i}$ is a polynomial in one variable with degree at least $2$ and $a_{i}\neq 0$ for every $1\leq i\leq m$. Such maps $H$ and their factors $H_{i}$ as in (1.1) will be referred to as generalised and simple Hénon maps respectively. Note that the degree of the first and second coordinates of $H$ are
[…]
respectively, and that the Jacobian determinant of $H$ is a constant and equals $a=a_{1}a_{2}\ldots a_{m}$.
Let ${\rm Aut}(\mathbb{C}^{2})$ denote the group of holomorphic automorphisms of $\mathbb{C}^{2}$. We identify all elements of ${\rm Aut}(\mathbb{C}^{2})$ that preserve the escaping set (or equivalently, the non-escaping set) associated with a generalised Hénon map $H$. Before proceeding, let us first recall the following definitions. The non-escaping set or the bounded Julia set associated with a generalised Hénon map $H$ is defined as $K^{+}=\{z\in\mathbb{C}^{2}:\|H^{n}(z)\|\text{ is bounded}\}$ and the escaping set is $U^{+}=\mathbb{C}^{2}\setminus K^{+}$. Furthermore, $H$ extends to $\mathbb{P}^{2}$ as a birational map with an indeterminacy at the point $I^{+}=[1:0:0]$ and a super-attracting fixed point at $I^{-}=[0:1:0]$. The escaping set $U^{+}$ corresponds to the intersection of the basin of attraction of $I^{-}=[0:1:0]$ with $\mathbb{C}^{2}$, and $I^{+}$ is the only accumulation point of the non-escaping set $K^{+}$ in $\mathbb{P}^{2}$. See [3], [13] for these details. The inverse of $H$,
$H^{-1}(x,y)=H_{1}^{-1}\circ\cdots\circ H_{m}^{-1}(x,y)$
has an indeterminacy at $I^{-}=[0:1:0]$ when extended to $\mathbb{P}^{2}$ and its non-escaping and escaping sets are defined as $K^{-}=\{z\in\mathbb{C}^{2}:\|H^{-n}(z)\|\text{ is bounded}\}\text{ and }U^{-}=\mathbb{C}^{2}\setminus K^{-}$ respectively. Furthermore, $U^{-}$ corresponds to the basin of attraction of $I^{+}=[1:0:0]$ of $H^{-1}$ when extended to $\mathbb{P}^{2}$ and $\overline{K^{-}}=K^{-}\cup I^{-}$.
The phenomenon of rigidity for generalised and simple maps Hénon has various manifestations. To place this work in context, it is useful to recall a few related results.
First, given a pair of generalised Hénon maps $H$ and $F$ that have the same non-escaping set, Lamy [16] showed that $H,F$ must be the same up to possibly different iterates. That is, there exist positive integers $m,n$ such that $H^{m}=F^{n}$. As a consequence, $H^{-1}$ and $F^{-1}$ will also have the same non-escaping and escaping sets.
Second, the notion of very rigid appears in the work of Dinh–Sibony [10] who showed that there exists a unique positive $dd^{c}$-closed $(1,1)$ current of mass $1$ on $K^{+}$. This extends results from [3], [13], and leverages the potential-theoretic properties of the Green’s function associated with $K^{+}$.
Third, [5] shows that if $f\in{\rm Aut}(\mathbb{C}^{2})$ preserves both $K^{\pm}$, then the following dichotomy holds: either $f$ is linear or one of $f$ or $f^{-1}$ must be a generalised Hénon map that commutes with $H$ up to a diagonal unimodular change of variables. As a consequence, the non-escaping sets of $f$ (or $f^{-1}$ as the case may be) and $H$ must be the same. The proof uses the very rigid property of $K^{\pm}$ alluded to above and the dependence of the Green’s functions on the Böttcher functions near the super-attracting fixed points $I^{\pm}$ in $\mathbb{P}^{2}$ constructed in [15]. Furthermore, by combining the above results, it is shown in [6] that if $f$ is an affine map that preserves the non-escaping set of a simple Hénon map $H$, then $f$ is conjugate to the linear map $(x,y)\mapsto(\eta x,\eta^{d}y)$ where $\eta$ is a $(d^{2}-1)$-root of unity, and $f\circ H^{2}=H^{2}\circ f$.
Next, Pal in [18] showed that two simple Hénon maps $H,F$ with the same degree that have biholomorphic escaping sets are related via affine automorphisms $A_{1}$ and $A_{2}$, i.e., $H=A_{1}\circ F\circ A_{2}$. This generalises an earlier result of Bonnot–Radu–Tanase [7].
Finally, and more recently, Cantat–Dujardin [9] show that if a pair of polynomial automorphisms $f,g$ in ${\rm Aut}(\mathbb{C}^{2})$ with dynamical degree $d\geq 2$ are conjugate by some $\varphi\in{\rm Aut}(\mathbb{C}^{2})$, then $\varphi$ must be a polynomial automorphism. This confirms a conjecture of Friedland–Milnor and can be viewed as a rigidity result for generalised Hénon maps. Related to this circle of ideas on conjugacy and motivated by a question of Favre [11], Bacher [1] shows that a pair of generalised Hénon maps are conjugate in ${\rm Aut}(\mathbb{C}^{2})$ precisely when the Kato surfaces associated to them near the super-attracting fixed point $I^{-}$ are biholomorphic.
The results in [6] and [18] build on the unpublished work of Bousch [8] that uses the description of an intermediate cover of $U^{+}$ given by Hubbard–Oberste-Vorth in [15] to compute the automorphism group of $U^{+}$ for simple quadratic Hénon maps. Our goal is to further generalise these results by identifying all automorphisms of $\mathbb{C}^{2}$ that preserve $U^{+}$ associated with generalised Hénon maps. Here is the main result:
Theorem 1.1.
Let $H$ be a generalised Hénon map of the form (1.1), i.e.,
$H(x,y)=H_{m}\circ\cdots\circ H_{1}(x,y)$
where $H_{i}(x,y)=(y,p_{i}(y)-a_{i}x)$, $p_{i}$ is a polynomial with degree $d_{i}\geq 2$ and $a_{i}\neq 0$ for every $1\leq i\leq m$. Then the non-escaping set $K^{+}$, or equivalently, the escaping set $U^{+}$, associated with $H$ is rigid under the action of ${\rm Aut}(\mathbb{C}^{2})$. In other words, any $f\in{\rm Aut}(\mathbb{C}^{2})$ that satisfies
$f(K^{+})=K^{+}\text{ or equivalently }f(U^{+})=U^{+},$
is of the form $f=L\circ H^{s}\text{ with }s\in\mathbb{Z},$ where $L$ is an affine map that preserves both $K^{\pm}$. In addition,
$\mathcal{L}=\{L:L\text{ is an affine map such that }L(K^{\pm})=K^{\pm}\}\leq\mathbb{Z}_{(d+d^{\prime})(d-1)}$
Here, the assumption that $f(K^{-})=K^{-}$, which was imposed in [5] is not required, and this is the main thrust of this theorem. The proof consists of several steps that are summarized as follows.
First, Theorem 2.1 in Section $2$ extends the construction of the intermediate cover of $U^{+}$ by Hubbard–Oberste-Vorth in [15] to the case of generalised Hénon maps. Next, let ${\rm Aut}_{1}(U^{+})$ be the subgroup of the holomorphic automorphism group ${\rm Aut}(U^{+})$ whose elements induce the identity map on the fundamental group $\pi_{1}(U^{+})$. Theorem 2.1 leads to a better understanding of the structure of ${\rm Aut}(U^{+})$. This is the content of Theorem 3.1, which shows that modulo a copy of $\mathbb{C}$, ${\rm Aut}(U^{+})$ is at most finite cyclic. The proof relies on adapting several ideas from [8]. Using this information about the structure of ${\rm Aut}(U^{+})$, Theorem 4.1 identifies all elements of ${\rm Aut}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ whose lifts to the intermediate cover of $U^{+}$ constructed above have a specific form. The final step is completed in Theorem 4.7 which extends the observations made in the proof of Theorem 4.1 and shows that ${\rm Aut}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ is a finite cyclic group whose order can be bounded above in terms of the degree of $H$. Similar results were observed earlier in [4].
Recall that the Green’s functions associated to the sets $K^{\pm}$ are defined as
[…]
They are non-negative, continuous, plurisubharmonic functions on $\mathbb{C}^{2}$ and satisfy the functorial property
(1.2) $G_{H}^{\pm}(H^{s}(z))=d^{\pm s}G_{H}^{\pm}(z)$
for every $s\in\mathbb{Z}$. Further, $K^{\pm}=\{G_{H}^{\pm}=0\}$ and
[…]
are $dd^{c}$-closed positive $(1,1)$ currents of mass 1 supported on $K_{H}^{+}$ and $K_{H}^{-}$, respectively. As indicated earlier, $\mu_{H}^{\pm}$ are precisely those currents that bestow on $K_{H}^{\pm}$ their very rigid property.
For every $c>0$, the sub-level set
$\Omega_{c}=\{z\in\mathbb{C}^{2}:G_{H}^{+}(z)<c\}$
is known to be a Short $\mathbb{C}^{2}$. Short $\mathbb{C}^{2}$’s are domains whose infinitesimal Kobayashi metric vanishes identically but at the same time admit a non-constant bounded plurisubharmonic function and which can be exhausted by an increasing union of domains biholomorphic to the ball. They were first observed and constructed by Fornæss in [12].
In [6] the analytical and topological structure of an intermediate cover of a punctured Short $\mathbb{C}^{2}$, i.e., $\Omega_{c}^{\prime}=\Omega_{c}\setminus K^{+}$, corresponding to a simple Hénon map was explored in a manner similar to [15], [7] and [8]. The philosophy behind this analysis is to understand the rigidity of an appropriate punctured neighbourhood of the super-attracting fixed point $I^{-}$ of $H$, in contrast to the rigidity of the Kato surfaces associated to the neighbourhood of $I^{-}$, raised by Favre in [11].
As a consequence, it is proved therein that the holomorphic automorphism group of certain punctured Short $\mathbb{C}^{2}$’s corresponding to a simple Hénon map can at most be finitely many copies of $\mathbb{C}$. Let ${\rm Aut}(\Omega_{c})$ and ${\rm Aut}(\Omega^{\prime}_{c})$ denote the holomorphic automorphism groups of $\Omega_{c}$ and $\Omega^{\prime}_{c}$ respectively. The subgroup of ${\rm Aut}(\Omega^{\prime}_{c})$ consisting of those elements that induce the identity map on the fundamental group $\uppi_{1}(\Omega^{\prime}_{c})$ will be denoted by ${\rm Aut}_{1}(\Omega^{\prime}_{c})$.
Theorem 2.1 has consequences for understanding both ${\rm Aut}(\Omega_{c})$ and ${\rm Aut}_{1}(\Omega^{\prime}_{c})$. First, we observe in Theorem 3.2 that covers of punctured Short $\mathbb{C}^{2}$’s obtained as sub-level sets of generalised Hénon maps can be constructed. This leads to Corollary 3.5 in which the automorphism group of such punctured Short $\mathbb{C}^{2}$’s can be identified in the same spirit as [6, Theorem 1.2]. Furthermore, we will use this representation to conclude the following:
Theorem 1.2.
Let $H$ be a generalised Hénon map. Then
${\rm Aut}_{1}(U^{+})={\rm Aut}_{1}(\Omega_{c}^{\prime})$
for every $c>0$.
By combining Theorem 1.1 and Theorem 1.2, we obtain
Corollary 1.3.
Let $H$ be a generalised Hénon map and $f\in{\rm Aut}(\Omega_{c})$ for some $c>0$. Then there exists an affine map $L\in{\rm Aut}(\mathbb{C}^{2})$ such that $L(K^{\pm})=K^{\pm}$ and $L(z)=f(z)$ on $\Omega_{c}$.
In other words,
Corollary 1.4.
Let $H$ be a generalised Hénon map. Then for every $c>0$, every element of ${\rm Aut}(\Omega_{c})$ is the restriction of an affine map that preserves the escaping set (or non-escaping set) of $H$ and ${\rm Aut}(\Omega_{c})$ is a finite cyclic group of order at most $\mathbb{Z}_{d_{0}(d-1)}$. Here, $d_{0}$ is an integer that depends on $d+d^{\prime}$.
Acknowledgements:
The first author is partially funded by the Mathematical Research Impact Centric Support (MTR/2023/001435) grant from the Anusandhan National Research Foundation (ANRF) of India.
2. An intermediate cover of the escaping set $U^{+}$
The goal of this section is to prove an analog of the Hubbard–Oberste-Vorth result on covering spaces of $U^{+}$ in [15] for generalized Hénon maps of the form (1.1). Here is the precise statement.
Theorem 2.1.
Associated with the subgroup $\mathbb{Z}$ of $\uppi_{1}(U^{+})$, there exists an open cover $\widehat{U}^{+}$ of $U^{+}$ that is biholomorphic to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. The generalised Hénon map $H$ admits a lift to $\widehat{U}^{+}$ and the lift is of the form
(2.1) $\widetilde{H}(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta),\zeta^{d}\right),$
where
$Q(\zeta)=\zeta^{d+d^{\prime}}+A_{d+d^{\prime}-1}\zeta^{d+d^{\prime}-1}+\ldots+A_{0}.$
Further,
  1. (i)
    The fundamental group $\uppi_{1}(U^{+})$ is isomorphic to $\mathbb{Z}\left[\frac{1}{d}\right]$.
  2. (ii)
    An element $\left[\frac{k}{d^{n}}\right]$ of $\uppi_{1}(U^{+})/\mathbb{Z}$ corresponds to the following deck transformation of $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$:
    (2.2) $\upgamma_{k/d^{n}}(z,\zeta)=\left(z+\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right),e^{2\pi ik/d^{n}}\cdot\zeta\right).$
This is a generalization of the results in [15, Section 8]. We will briefly revisit the steps, mostly following [17, Chapter 7]. Note that
$H=H_{m}\circ H_{m-1}\circ\cdots\circ H_{1}\text{ with }H_{i}(x,y)=(y,p_{i}(y)-a_{i}x),$
where $p_{i}$ are monic polynomials of degree $d_{i}$, $1\leq i\leq m$. Let us first recall from [3] and [15], a few properties of the dynamics of a Hénon map of the above form at infinity. Consider the filtration of $\mathbb{C}^{2}$ by
$V_{R}=\left\{|x|,|y|\leq R\right\},V^{+}_{R}=\left\{|y|\geq\max\{|x|,R\}\right\},\;V^{-}_{R}=\left\{|x|\geq\max\{|y|,R\}\right\}.$
Then, there exists a large $R_{H}>1$ such that for every $R\geq R_{H}$
(2.3) $H(V^{+}_{R})\subset V^{+}_{R}\text{ and }H^{-1}(V^{-}_{R})\subset V^{-}_{R}.$
Further,
[…]
Thus, $K^{\pm}\subset\textsf{int}\left(V_{R}\cup V_{R}^{\mp}\right)$ and $\mathbb{C}^{2}\setminus\textsf{int}\left(V_{R}\cup V_{R}^{\mp}\right)\subset U^{\pm}$. The value $R_{H}>1$ corresponding to the map $H$ (1.1) will be referred to as the radius of filtration of $H$ and we will write $V^{\pm}$ in place of $V_{R_{H}}^{\pm}$ unless stated otherwise.
The proof of Theorem 2.1 involves several steps and is divided into subsections accordingly.
2.1. Construction of the cover $\widehat{U}^{+}$
Let us first recall from [5] some results on the construction of the Böttcher coordinate function for Hénon maps. Let $(x_{n},y_{n})=H^{n}(x,y)$ for $(x,y)\in V^{+}$. Define
$y_{n+1}=y^{d}_{n}\Big(1+\left({q(x_{n},y_{n})}/{y^{d}_{n}}\right)\Big)\sim y^{d}_{n}$
for $(x,y)\in V^{+}.$ Now by Proposition 2.1 from [5]
$\phi(x,y)=y\lim_{n\rightarrow\infty}\prod_{j=0}^{n-1}\left(1+\frac{q(x_{j},y_{j})}{y^{d}_{j}}\right)^{1/d^{j+1}}$
is a well defined holomorphic function on $V^{+}$ as the limit is uniform on compact subsets of $V^{+}$. In addition, $\phi\circ H=\phi^{d}$ and $G_{H}^{+}=\log|\phi|$ on $V^{+}$ with
(2.5) $\phi(x,y)=ye^{\gamma(x,y)}\text{ and }\gamma(x,y)\to 0\text{ as }y\to\infty.$
Note that $\omega=\frac{d\phi}{\phi}$ is a closed 1-form in $V^{+}$ and $H^{*}\omega=d\omega$ where $d\geq 2$ is the degree of $H$. Hence, as in Proposition 7.3.2 in [17], $\omega$ extends to a closed 1-form on $U^{+}$ by (2.4). For a closed curve $C$ in $U^{+}$ define
$\alpha(C)=\frac{1}{2\pi i}\int_{C}\omega.$
Also, Proposition 7.3.3 in [17], which is stated for simple Hénon maps, generalises as:
Proposition 2.2.
For a closed curve $C$ in $U^{+}$, the following holds:
  • (i)
    $\alpha(H(C))=d\alpha(C)$, where $H(C)$ is the image of the curve $C$ under $H$.
  • (ii)
    $\alpha(C)\in\mathbb{Z}\left[\frac{1}{d}\right]$ where $\mathbb{Z}\left[\frac{1}{d}\right]=\left\{\frac{k}{d^{n}}:k,n\in\mathbb{Z}\right\}.$
  • (iii)
    $\alpha:\displaystyle\uppi_{1}(U^{+})\rightarrow\mathbb{Z}\left[\frac{1}{d}\right]$ is a group isomorphism.
Further, as in Chapter $7$ of [17], let $\mathcal{H}$ be the subgroup of $\uppi_{1}(U^{+})\simeq\mathbb{Z}[\frac{1}{d}]$ that corresponds to $\mathbb{Z}$, i.e., $\alpha(\mathcal{H})=\mathbb{Z}$. Fix a point $z_{0}\in V^{+}$ and consider all the pairs $(z,C)$ of points $z\in U^{+}$, where $C$ is a path in $U^{+}$ with initial point $z_{0}$ and end point $z$. Define an equivalence relation
$(z,C)$ $\sim(z^{\prime},C^{\prime})\text{ if and only if }z=z^{\prime}\text{ and }CC^{\prime-1}\in\mathcal{H}$.
The covering space $\widehat{U}^{+}$ is defined as
[…]
The projection map $\widehat{\pi}:\widehat{U}^{+}\to U^{+}$ is defined as $\widehat{\pi}([z,C])=z.$ Equip $\widehat{U}^{+}$ with the pull-back complex structure so that the projection map $\widehat{\pi}$ becomes holomorphic. Note that $\widehat{\pi}$ is a covering map as it is a local homeomorphism, and the fundamental group of $\widehat{U}^{+}$ is $\mathbb{Z}$.
Consider the holomorphic map $\hat{\phi}$ defined on $\widehat{U}^{+}$ as
$\hat{\phi}(p)=\phi(z_{0})\exp\left(\int_{C}\omega\right),$
where $\phi$ is the Böttcher coordinate (2.5) defined on $V_{R}^{+}$. Let $\widehat{H}$ be the lift of $H$ to $\widehat{U}^{+}$ defined as
$\widehat{H}([z,C]=[H(z),lH(C)],$
where $l$ is a path in $V^{+}$ connecting $z_{0}$ and $H(z_{0})$. Now Propositions 7.3.5 and 7.3.6 from [17] collectively generalise as:
Proposition 2.3.
[…]
  • (i)
    for $p\in\widehat{V}^{+}$, $\hat{\phi}(p)=\phi(\widehat{\pi}(p))$ and $\widehat{\pi}$ is a biholomorphism between $\widehat{V}^{+}$ and $V^{+}$.
  • (ii)
    for $p,p^{\prime}\in\widehat{U}^{+}$ if $\hat{\phi}(p)=\hat{\phi}(p^{\prime})$ and $\widehat{\pi}(p)=\widehat{\pi}(p^{\prime})$, then $p=p^{\prime}$.
  • (iii)
    the functional equation $\hat{\phi}(\widehat{H}(p))=(\hat{\phi}(p))^{d}$ holds.
The proof is exactly the same as in [17]. Also $\widehat{H}(\widehat{V}^{+})\subset\widehat{V}^{+}.$ The next step is to construct
2.2. The form of the map $\widehat{H}$ on an appropriate subdomain
Recall that $R_{H}>1$ is the radius of filtration of the map $H$. For every $M>1$ define the sets
$W_{M}^{+}=\left\{(x,y)\in\mathbb{C}^{2}:|y|>M\max\{|x|,R_{H}\}\right\}$
and
$\widetilde{W}_{M}^{+}=\left\{(x,y)\in\mathbb{C}^{2}:|\phi(x,y)|>M\max\{|x|,R_{H}\}\right\}.$
Note that by (2.5), there exists $M_{0}>1$ such that for all $M>M_{0}$, $\widetilde{W}_{M}^{+}\subset V^{+}$. Hence, consider the map $E_{1}:\widetilde{W}_{M}^{+}\to W_{M}^{+}$ given by
$E_{1}(x,y)=(x,\phi(x,y)).$
Then $E_{1}(\widetilde{W}_{M}^{+})=W_{M}^{+}$ and $E_{1}$ is a biholomorphism whose inverse is
$E_{1}^{-1}(x,y)=(x,\lambda(x,y))$
for some holomorphic $\lambda:W_{M}^{+}\to\mathbb{C}$.
Lemma 2.4.
For $\epsilon>0$ sufficiently small, there exists $M_{\epsilon}>M_{0}$ such that
  • (i)
    for $(x,y)\in\widetilde{W}_{M_{\epsilon}}^{+}$
[…]
  • (ii)
    for $(x,y)\in W_{M_{\epsilon}}^{+}$
[…]
Proof.
From (2.5) there exists $M_{1}\geq 0$ such that
$1-\epsilon\leq\left|\frac{\phi(x,y)}{y}\right|\leq 1+\epsilon\text{ and }1-\epsilon\leq\left|\frac{\lambda(\tilde{x},\tilde{y})}{\tilde{y}}\right|\leq 1+\epsilon$
for $(x,y)\in\widetilde{W}_{M_{1}}^{+}$ and $(\tilde{x},\tilde{y})\in W_{M_{1}}^{+}$ respectively. Let $D^{2}(z;1)$ denote the polydisk of polyradius $1$ centered at $z$. Now choose $M_{\epsilon}>M_{1}$ such that $D^{2}(z;1)\subset\widetilde{W}_{M_{1}}^{+}$ and $D^{2}(\tilde{z};1)\subset W_{M_{1}}^{+}$ for every $z\in\widetilde{W}_{M_{\epsilon}}^{+}$ and $\tilde{z}\in W_{M_{\epsilon}}^{+}$ respectively. By the Cauchy estimates,
$\left|\frac{\partial\phi}{\partial y}(x,y)\right|\leq 1+\epsilon\text{ and }\left|\frac{\partial\lambda}{\partial y}(\tilde{x},\tilde{y})\right|\leq 1+\epsilon,$
for $(x,y)\in\widetilde{W}_{M_{\epsilon}}^{+}$ and $(\tilde{x},\tilde{y})\in W_{M_{\epsilon}}^{+}$ respectively. Since $E_{1}(\widetilde{W}_{M_{\epsilon}}^{+})=W_{M_{\epsilon}}^{+}$ and $\lambda(x,\phi(x,y))=y$,
$\left|\frac{\partial\phi}{\partial y}(x,y)\right|=\left|\frac{\partial\lambda}{\partial y}\left(x,\phi(x,y)\right)\right|^{-1}\geq(1+\epsilon)^{-1}\geq 1-\epsilon.$
[…]
Remark 2.5.
By (2.5), $e^{\gamma(x,y)}\to 1$ as $y\to\infty$ on $V^{+}$. Since
$\frac{\partial\phi}{\partial y}(x,y)=e^{\gamma(x,y)}+\phi(x,y)\gamma_{y}(x,y)$
it follows from Lemma 2.4 that for $\epsilon>0$ sufficiently small,
  • (i)
[…]
  • (ii)
[…]
For the next part of the computations in this subsection, we will fix the constant $M>M_{\epsilon}$ where $0<\epsilon<1/2$ is sufficiently small such that $H(\widetilde{W}_{M}^{+})\subset\textsf{int}\left(\widetilde{W}_{M}^{+}\right)$. Consider the map $E_{2}$ on $W_{M}^{+}$ defined as
$E_{2}(x,y)=(\psi(x,y),y)\text{ where }\psi(x,y)=y\int_{0}^{x}\frac{\partial\lambda}{\partial y}(t,y)dt.$
Note that $\psi$ is well-defined on $W_{M}^{+}$ by Lemma 2.4. Next, we claim that $E_{2}$ is injective on $W_{M}^{+}$. To see this, suppose that $E_{2}(x_{1},y_{1})=E_{2}(x_{2},y_{2})$ on $W_{M}^{+}$. Then $y_{1}=y_{2}=y\neq 0$ and
$y\int_{x_{1}}^{x_{2}}\frac{\partial\lambda}{\partial y}(t,y)dt=0,\text{ i.e., }\int_{x_{1}}^{x_{2}}\frac{\partial\lambda}{\partial y}(t,y)dt=0,$
where $x_{1},x_{2}$ belong to the disc $D\left(0,|y|/M\right)\subset\mathbb{C}$. In particular, by Remark 2.5 (ii) and the choice of $\epsilon>0$,
(2.6) $|x_{1}-x_{2}|\leq\left|\int_{x_{1}}^{x_{2}}\left(\frac{\partial\lambda}{\partial y}(t,y)-1\right)dt\right|\leq\frac{1}{2}|x_{1}-x_{2}|.$
Hence $x_{1}=x_{2}$ and this shows the injectivity of $E_{2}$ on $W_{M}^{+}$. Now consider the map $\Psi=E_{2}\circ E_{1}$ on $\widetilde{W}_{M}^{+}$. Again, as $\lambda(x,\phi(x,y))=y$ for every $(x,y)\in\widetilde{W}_{M}^{+}$,
(2.7) $\text{Det }D\Psi(x,y)=\phi(x,y)\frac{\partial\phi}{\partial y}(x,y)\frac{\partial\lambda}{\partial y}\left(x,\phi(x,y)\right)=\phi(x,y)\neq 0.$
Let $\Omega_{M}^{+}=\Psi(\widetilde{W}_{M}^{+})$ which is a domain in $\mathbb{C}^{2}$. Then $\Psi$ is a biholomorphism from $\widetilde{W}_{M}^{+}$ onto $\Omega_{M}^{+}$.
Lemma 2.6.
Let $\widetilde{H^{\prime}}$ be the lift of the map $H$ from $\widetilde{W}_{M}^{+}$ to $\Omega_{M}^{+}$, i.e., $\widetilde{H^{\prime}}=\Psi\circ H\circ\Psi^{-1}$. Then there exists a holomorphic map $\tilde{Q}$ such that
$\widetilde{H^{\prime}}(z,\zeta)=\left(\frac{a}{d}z+\tilde{Q}(\zeta),\zeta^{d}\right).$
Also for $|\zeta|>MR_{H}$, $\tilde{Q}(\zeta)=\zeta^{d+d^{\prime}}+A_{d+d^{\prime}-1}\zeta^{d+d^{\prime}-1}+\cdots+A_{0}+\sum_{i=1}^{\infty}A_{-i}\zeta^{-i}$, where $d,d^{\prime}$ are as defined before.
Proof.
Let $h_{1}(z,\zeta)=\pi_{1}\circ\widetilde{H^{\prime}}(z,\zeta)$ and $h_{2}(z,\zeta)=\pi_{2}\circ\widetilde{H^{\prime}}(z,\zeta)$ be the components of $\widetilde{H^{\prime}}$. From the properties of the Böttcher coordinate $\phi$ as discussed in subsection 2.1, it follows that for $(x,y)\in\widetilde{W}_{M}^{+}$
$h_{2}(\Psi(x,y))=\phi\circ H(x,y)=\phi(x,y)^{d},\text{ i.e., }h_{2}(z,\zeta)=\zeta^{d}$
for $(z,\zeta)\in\Omega_{M}^{+}$. In particular, $\frac{\partial h_{2}}{\partial\zeta}(z,\zeta)=d\zeta^{d-1}$ and $\frac{\partial h_{2}}{\partial z}(z,\zeta)=0.$ Thus
$\text{Det }D\widetilde{H^{\prime}}(z,\zeta)=d\zeta^{d-1}\frac{\partial h_{1}}{\partial z}(z,\zeta).$
Now on $\widetilde{W}_{M}^{+}$, $\widetilde{H^{\prime}}\circ\Psi=\Psi\circ H$. Hence from (2.7)
$\text{Det }D\widetilde{H^{\prime}}(\Psi(x,y))=\text{Det }D\widetilde{H^{\prime}}(\psi(x,\phi(x,y)),\phi(x,y))=a\phi(x,y)^{d-1},$
i.e., $\text{Det }D\widetilde{H^{\prime}}(z,\zeta)=a\zeta^{d-1}=d\zeta^{d-1}\frac{\partial h_{1}}{\partial z}(z,\zeta).$ So for every $(z,\zeta)\in\Omega_{M}^{+}$
$\frac{\partial h_{1}}{\partial z}(z,\zeta)=\frac{a}{d}\text{ or }h_{1}(z,\zeta)=\frac{a}{d}z+\tilde{Q}(\zeta)$
where $\tilde{Q}$ is holomorphic on $\Omega_{M}^{+}$. To prove that $\tilde{Q}$ has the desired form, it suffices to show $Q(\zeta)/\zeta^{d+d^{\prime}}\to 1$ as $|\zeta|\to\infty$. Since $E_{2}^{-1}(0,\zeta)=(0,\zeta)$,
$\tilde{Q}(\zeta)=h_{1}(0,\zeta)$ $=\psi\circ H\circ E_{1}^{-1}(0,\zeta)=\psi(H(0,\lambda(0,\zeta)))=\psi(P_{1}(\lambda(0,\zeta)),\phi\circ H(0,\lambda(0,\zeta)))$
$=\psi\left(P_{1}(\lambda(0,\zeta)),(\phi(0,\lambda(0,\zeta))^{d}\right)=\psi\left(P_{1}(\lambda(0,\zeta)),\zeta^{d}\right)$
where $P_{1}(x)=\pi_{1}\circ H(0,x)$, a monic polynomial of degree $d^{\prime}=dd_{m}^{-1}$. Thus
$\tilde{Q}(\zeta)=\psi\left(P_{1}\left(\lambda(0,\zeta)\right),\zeta^{d}\right)=\zeta^{d}\int_{0}^{P_{1}(\lambda(0,\zeta))}\frac{\partial\lambda}{\partial y}\left(t,\zeta^{d}\right)dt.$
Now, from Remark 2.5 (ii) it follows that
$\frac{P_{1}(\lambda(0,\zeta))}{\zeta^{d^{\prime}}}\to 1$
as $\zeta\to\infty$. Since $2d^{\prime}\leq d$, for $|\zeta|$ sufficiently large, $(t,\zeta^{d})\in W_{M}^{+}$ for every $0\leq|t|\leq|P_{1}(\lambda(0,\zeta))|$. In particular, from Remark 2.5 (ii) again, it follows that as $\zeta\to\infty$,
$\frac{\partial\lambda}{\partial y}\left(t,\zeta^{d}\right)\to 1\text{ whenever }|t|\leq|P_{1}(\lambda(0,\zeta))|.$
Thus, as $\zeta\to\infty$,
$\left|\frac{\tilde{Q}(\zeta)}{\zeta^{d+d^{\prime}}}-\frac{P_{1}(\lambda(0,\zeta))}{\zeta^{d^{\prime}}}\right|\to 0,$
and consequently $\frac{\tilde{Q}(\zeta)}{\zeta^{d+d^{\prime}}}\to 1$. This completes the proof. ∎
The Laurent expansion of $\tilde{Q}$ can be written as $\tilde{Q}(\zeta)=Q(\zeta)+Q^{-}(\zeta)$ where $Q$ is the principal part and $Q^{-}(\zeta)=\sum_{i=1}^{\infty}A_{-i}\zeta^{-i}$. Since $|\zeta|>MR_{H}$ for every $(z,\zeta)\in\Omega_{M}^{+}$ and $Q^{-}(\zeta)$ is defined there, the functions
$R_{n}(\zeta)=\sum_{i=0}^{n}\left(\frac{d}{a}\right)^{i+1}Q^{-}\left(\zeta^{d^{i}}\right)$
are well-defined for such $\zeta$. Then for every $|\zeta|>MR_{H}$, there exists $n_{\zeta}\geq 1$ and a constant $C>0$ such that
$|R_{n}(\zeta)-R_{n-1}(\zeta)|\leq C\left|\frac{d}{a}\right|^{n}|\zeta|^{-d^{n}},$
whenever $n\geq n_{\zeta}$. Hence $\{R_{n}\}$ converges locally uniformly to a holomorphic function, say $R$ on $\Omega_{M}^{+}$. Consider the injective map $E_{3}$ on $\Omega_{M}^{+}$ as $E_{3}(x,y)=(x-R(y),y)$. Let $\widetilde{\Psi}=E_{3}\circ\Psi$ and
$\widetilde{\Omega}_{M}^{+}=E_{3}(\Omega_{M}^{+})=\widetilde{\Psi}(\widetilde{W}_{M}^{+}).$
Note that for $(z,\zeta)\in\widetilde{\Omega}_{M}^{+}$,
$\left(\frac{a}{d}\right)R(\zeta)-R(\zeta^{d})=Q^{-}(\zeta)$
and hence
$\widetilde{\Psi}\circ H\circ\widetilde{\Psi}^{-1}(z,\zeta)=E_{3}\circ\widetilde{H^{\prime}}\circ E_{3}^{-1}(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta),\zeta^{d}\right),$
which proves the following
Proposition 2.7.
Let $\widetilde{H}=\widetilde{\Psi}\circ H\circ\widetilde{\Psi}^{-1}$ on $\widetilde{\Omega}_{M}^{+}$. Then for $(z,\zeta)\in\widetilde{\Omega}_{M}^{+}$
$\widetilde{H}(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta),\zeta^{d}\right),$
where $Q(\zeta)=\zeta^{d+d^{\prime}}+A_{d+d^{\prime}-1}\zeta^{d+d^{\prime}-1}+\cdots+A_{0}$.
2.3. $\widehat{U}^{+}$ is biholomorphic to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$
Recall from Proposition 2.3 that $\widehat{\pi}$ restricted to $\widehat{V}^{+}$ is a biholomorphism onto its image. Let $\widehat{\pi}(\widehat{W}_{M}^{+})=\widetilde{W}_{M}^{+}$ where $\widehat{W}_{M}^{+}\subset\widehat{V}^{+}\subset\widehat{U}^{+}$ and $M>0$ is sufficiently large as fixed in the previous subsection. As $\widehat{H}$ is the lift of $H$ to $\widehat{U}^{+}$,
[…]
for every $n\geq 1$. By Proposition 2.7, $H(\widetilde{W}_{M}^{+})\subset\widetilde{W}_{M}^{+}$on $\widetilde{W}_{M}^{+}$ and hence
[…]
on $\widetilde{W}_{M}^{+}$ and for every $n\geq 1$. Thus, we observe the following:
Proposition 2.8.
Both $\left\{\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)\right\}$ and $\left\{\widetilde{H}^{-n}\left(\widetilde{\Omega}_{M}^{+}\right)\right\}$ are increasing sequences of domains with
$\widehat{U}^{+}=\bigcup_{n=0}^{\infty}\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)\text{ and }\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)=\bigcup_{n=0}^{\infty}\widetilde{H}^{-n}\left(\widetilde{\Omega}_{M}^{+}\right).$
Proof.
Note that $I^{-}=[0:1:0]$ is a super-attracting fixed point of $H$ when extended to $\mathbb{P}^{2}$ with $U^{+}$ as its basin of attraction. Furthermore, $\widetilde{W}_{M}^{+}$ corresponds to a neighbourhood of $I^{-}$ in $\mathbb{P}^{2}$. Thus, we have
$U^{+}=\bigcup_{n=0}^{\infty}H^{-n}\left(\widetilde{W}_{M}^{+}\right)$
and $\left\{{H}^{-n}\left(\widetilde{W}_{M}^{+}\right)\right\}$ is an increasing sequence. By (2.8),
$H^{n}\circ\widehat{\pi}\circ\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)=\widehat{\pi}\left(\widehat{W}_{M}^{+}\right)=\widetilde{W}_{M}^{+}$
for every $n\geq 1$. As $H$ is an automorphism of $\mathbb{C}^{2}$, $\widehat{\pi}\circ\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)={H}^{-n}\left(\widetilde{W}_{M}^{+}\right)$. Since $\left\{{H}^{-n}\left(\widetilde{W}_{M}^{+}\right)\right\}$ is an increasing sequence, $\left\{\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)\right\}$ is also an increasing sequence of domains with $\widehat{U}^{+}=\bigcup_{n=0}^{\infty}\widehat{H}^{-n}\left(\widehat{W}_{M}^{+}\right)$. Consider $\mathbb{C}^{2}$ with coordinates $(z,\zeta)$ and the maps
$P_{1}(z,\zeta)=(z,\zeta^{d})\text{ and }P_{2}(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta),\zeta\right).$
Note that $\widetilde{H}=P_{1}\circ P_{2}$ on $\widetilde{W}_{M}^{+}$. Since $P_{1}$ is a proper self–map pf $\mathbb{C}^{2}$ and $P_{2}$ is an automorphism of $\mathbb{C}^{2}$, $\widetilde{H}$ extends to a proper map on $\mathbb{C}^{2}$. Further, as $H\left(\widetilde{W}_{M}^{+}\right)\subset\textsf{int}\left(\widetilde{W}_{M}^{+}\right)$, it follows from (2.9) that $\widetilde{H}\left(\widetilde{\Omega}_{M}^{+}\right)\subset\textsf{int}\left(\widetilde{\Omega}_{M}^{+}\right)$. Inductively, $\left\{\widetilde{H}^{-n}\left(\widetilde{\Omega}_{M}^{+}\right)\right\}$ is an increasing sequence. Since $|\zeta|>MR_{H}$ in $\widetilde{\Omega}_{M}^{+}$, $\widetilde{H}^{-n}\left(\widetilde{\Omega}_{M}^{+}\right)\subset\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ for every $n\geq 1$. Hence $\bigcup_{n=0}^{\infty}\widetilde{H}^{-n}\left(\widetilde{\Omega}_{M}^{+}\right)\subseteq\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. To complete the reasoning, we first observe the following:
Lemma 2.9.
There exists $\widetilde{M}>MR_{H}$ sufficiently large and $0<t<1$ such that
$S_{\widetilde{M},t}=\left\{(z,\zeta)\in\mathbb{C}^{2}:|z|<t|\zeta|^{2}\text{ and }|\zeta|\geq\widetilde{M}\right\}\subset\widetilde{\Omega}_{M}^{+}.$
[…]
Figure 1. Behaviour of $\widetilde{\Psi}$ on $\widetilde{W}_{M}^{+}$ and the image of $\widetilde{\Psi}(0,y)$.
Proof.
Recall the function $R(\zeta)$ from the proof of Proposition 2.7. Choose $\widetilde{M}>2MR_{H}$ sufficiently large, such that for $|\zeta|>\widetilde{M}$, $|R(\zeta)|<\frac{1}{4M}|\zeta|^{2}$. Fix $t=\frac{1}{4M}$. Then for $(z,\zeta)\in S_{\widetilde{M},t}$
$\left|z+R(\zeta)\right|<\frac{1}{2M}|\zeta|^{2}\text{ or }2M\left|\frac{z+R(\zeta)}{\zeta}\right|<|\zeta|.$
Hence $\left(\frac{z+R(\zeta)}{\zeta},\zeta\right)\in W_{2M}^{+}\subset W_{M}^{+}$. Recall the map $E_{2}$ from subsection 2.2 that is defined on $W_{M}^{+}$. In particular, $E_{2}(x,y)=\mathrm{P}\circ E(x,y)$, where
$E(x,y)=\left(\int_{0}^{x}\frac{\partial\lambda}{\partial y}(t,y)dt,y\right)\text{ and }\mathrm{P}(x,y)=(xy,y).$
Note that $\mathrm{P}$ is defined on $\mathbb{C}^{2}$ and $E$ on $W_{M}^{+}$. Also, $\mathrm{P}$ is injective on $\mathbb{C}\times\mathbb{C}^{*}$ and for a fixed $|y|>MR_{H}$, the slice $D\left(0,\frac{|y|}{M}\right)\times\{y\}\subset W_{M}^{+}$. So consider the function
$h_{y}(x)=\pi_{1}\circ E(x,y)$
that is well-defined on $D\left(0,\frac{|y|}{M}\right)$. By Remark 2.5(ii) and by a similar argument as in (2.6),
$|h_{y}(x)-x|\leq\epsilon|x|<\frac{\epsilon}{M}|y|\text{ and }h_{y}(0)=0,$
where $0<\epsilon<\frac{1}{2}$ as fixed there. Furthermore, $\overline{W_{M}^{+}}\subset V^{+}$ and $h_{y}$ extends continuously to $\partial D\left(0,\frac{|y|}{M}\right)$. Hence $h_{y}$ maps the boundary $\partial D(0,\frac{|y|}{M})$ into the annulus centered at the origin with inner radius $\frac{1-\epsilon}{M}|y|$ and outer radius $\frac{1+\epsilon}{M}|y|$.
We now claim that the entire disc $D\left(0,\frac{1-\epsilon}{M}|y|\right)$ is contained in $h_{y}\left(D\left(0,\frac{|y|}{M}\right)\right)$. To see this, note that there is a neighbourhood of the origin that is contained in the image $h_{y}\left(D\left(0,\frac{|y|}{M}\right)\right)$ as $h_{y}$ is a non-constant holomorphic map that fixes the origin. Let $\varrho>0$ be the largest disc centered at the origin that is contained in $h_{y}\left(D\left(0,\frac{|y|}{M}\right)\right)$. Suppose that $\varrho<\frac{1-\epsilon}{M}|y|$. There must exist a $b_{0}$ with $|b_{0}|=\varrho$ such that $b_{0}$ lies in the boundary of the image $h_{y}\left(D\left(0,\frac{|y|}{M}\right)\right)$. Choose $b_{n}\in h_{y}\left(D\left(0,\frac{|y|}{M}\right)\right)$ and pre-images $c_{n}\in h_{y}^{-1}(b_{n})\cap D(0,\frac{|y|}{M})$. Any limit point $c_{0}$ of the $c_{n}$’s must lie on the boundary $\partial D(0,\frac{|y|}{M})$ by the open mapping theorem. This implies that $|h_{y}(c_{0})|=|b_{0}|=\varrho<\frac{1-\epsilon}{M}|y|$ and this contradicts the fact that
$\frac{1-\epsilon}{M}|y|\leq|h_{y}(c_{0})|\leq\frac{1+\epsilon}{M}|y|.$
Since $\epsilon<\frac{1}{2}$, it follows from the above that $D\left(0,\frac{|y|}{2M}\right)\times\{y\}\subset E(W_{M}^{+})$ for every $|y|>MR_{H}$. Thus $E^{-1}\left(W_{2M}^{+}\right)\subset W_{M}^{+}$. In particular, if $(z,\zeta)\in S_{\widetilde{M},t}$, where $\widetilde{M}$ and $t$ are as fixed above, then
$E^{-1}\left(\frac{z+R(\zeta)}{\zeta},\zeta\right)\in W_{M}^{+}\text{ or }(z,\zeta)\in E_{3}\circ E_{2}(W_{M}^{+})=\widetilde{\Omega}_{M}^{+}.\qed$
Now, note that $\widetilde{H}$ is defined on $\mathbb{C}^{2}$. So for every $(z,\zeta)\in\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right).$
$\widetilde{H}^{n}(z,\zeta)$ $=\left(\left(\frac{a}{d}\right)^{n}z+\left(\frac{a}{d}\right)^{n-1}Q(\zeta)+\ldots+\frac{a}{d}Q\left(\zeta^{d^{n-2}}\right)+Q\left(\zeta^{d^{n-1}}\right),\zeta^{d^{n}}\right),$
for every $n\geq 1$. Let $h_{n}^{z}(\zeta)=\frac{\pi_{1}\circ\widetilde{H}^{n}(z,\zeta)}{\zeta^{2d^{n}}}$.
Claim: For every $z\in\mathbb{C}$ and $|\zeta|>1$, $\lim_{n\to\infty}h_{n}^{z}(\zeta)\to 0\text{ as }n\to\infty.$
Note that as $Q(\zeta)$ is a monic polynomial of degree $\zeta^{d+d^{\prime}}$, where $d^{\prime}=dd_{m}^{-1}$, there exists a constant $C_{1}>0$ such that
$\left|\frac{Q(\zeta)}{\zeta^{2d}}\right|<\left|\frac{1}{\zeta^{d-d^{\prime}}}\right|C_{1}$
for every $|\zeta|>1$. Now fix a point $(z,\zeta)\in\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. Then for every $n\geq 1$
(2.10) $h_{n+1}^{z}(\zeta)=\frac{a}{d}\frac{h_{n}^{z}(\zeta)}{\zeta^{2d^{n}(d-1)}}+\frac{Q(\zeta^{d^{n}})}{\left(\zeta^{d^{n}}\right)^{2d}}.$
In particular, for sufficiently large $n\geq 1$, $\left|\frac{a}{d}\zeta^{-2d^{n}(d-1)}\right|<1$, and
[…]
Thus $\{h_{n}^{z}(\zeta)\}$ is a bounded sequence, and from (2.10) there exist a constant $C>0$ such that
$\left|h_{n}^{z}(\zeta)\right|<\left|\frac{C}{\zeta^{d^{n}}}\right|,$
which proves the claim. Hence for sufficiently large $n$, $\widetilde{H}^{n}(z,\zeta)\in S_{\widetilde{M},t}\subset\widetilde{\Omega}_{M}^{+}$, where $\widetilde{M}$ and $t$ are as chosen in the proof of Lemma 2.9. This completes the proof of Proposition 2.8. ∎
Finally, we conclude this subsection with the following result.
Theorem 2.10.
$\widehat{U}^{+}$ is biholomorphic to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. Consequently, the universal cover of $U^{+}$ is $\mathbb{C}\times\mathbb{D}$.
Proof.
Note $\widetilde{\Psi}\circ\widehat{\pi}$ is a biholomorphism from $\widehat{W}_{M}^{+}$ to $\widetilde{\Omega}_{M}^{+}$ with
$\pi_{2}\circ\widetilde{\Psi}\circ\widehat{\pi}(p)=\hat{\phi}(p),$
where $\hat{\phi}$ is as defined in subsection 2.1 on $\widehat{U}^{+}$. Let $\hat{\psi}(p)=\pi_{1}\circ\widetilde{\Psi}\circ\widehat{\pi}(p).$ Then by Proposition 2.7, for $p\in\widehat{W}_{M}^{+}$
$\left(\hat{\psi}\circ\widehat{H}(p),\hat{\phi}\circ\widehat{H}(p)\right)=\widetilde{\Psi}\circ\widehat{\pi}\circ\widehat{H}(p)=\widetilde{H}\circ\widetilde{\Psi}\circ\widehat{\pi}(p)=\left(\frac{a}{d}\hat{\psi}(p)+Q\left(\hat{\phi}(p)\right),\hat{\phi}(p)^{d}\right).$
For $p\in\widehat{H}^{-1}(\widehat{W}_{M}^{+})$ define
$\hat{\psi}(p)=\frac{d}{a}\left(\hat{\psi}\left(\widehat{H}(p)\right)-Q\left(\hat{\phi}(p)\right)\right).$
Hence $\widetilde{\Psi}\circ\widehat{\pi}$ extends to $\widehat{H}^{-1}(\widehat{W}_{M}^{+})$ onto $\widetilde{H}^{-1}(\widetilde{\Omega}_{M}^{+})$. Now for $p,q\in\widehat{H}^{-1}(\widehat{W}_{M}^{+})$ if
$\widetilde{\Psi}\circ\widehat{\pi}(p)=\widetilde{\Psi}\circ\widehat{\pi}(q)$
then $\hat{\phi}(p)=\hat{\phi}(q)$ and
$\widetilde{H}\circ\widetilde{\Psi}\circ\widehat{\pi}(p)=\widetilde{\Psi}\circ H\circ\widehat{\pi}(p)=\widetilde{\Psi}\circ H\circ\widehat{\pi}(q)=\widetilde{H}\circ\widehat{\Psi}\circ\widetilde{\pi}(q).$
As $\widetilde{\Psi}$ is injective on $\widetilde{W}_{M}^{+}$, it follows that
[…]
Hence by Proposition 2.3 we have $p=q$. Furthermore, from Proposition 2.8, along with an induction argument it follows that $\widetilde{\Psi}\circ\widehat{\pi}$ extends to a biholomorphism from $\widehat{U}^{+}$ to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. The second part is immediate as the universal cover of $\mathbb{D}\setminus\{0\}$ is $\mathbb{D}$. ∎
2.4. Deck transformations of $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$
Note $\widetilde{\Psi}\circ\widehat{\pi}$ is a biholomorphism between $\widehat{U}^{+}$ and $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. Then $\widehat{\Pi}=\widehat{\pi}\circ\left(\widetilde{\Psi}\circ\widehat{\pi}\right)^{-1}$ is the covering map. Since $\widetilde{\Psi}$ and $\widehat{\pi}$ restricted to $\widetilde{W}_{M}^{+}$ and $\widehat{W}_{M}^{+}$ respectively are biholomorphisms, it follows that $\widehat{\Pi}=\widetilde{\Psi}^{-1}$ on $\widetilde{\Omega}_{M}^{+}$. Now by Proposition 2.7, $\widetilde{H}$ is the lift of the map $H$, i.e., $H=\widetilde{\Psi}\circ\widetilde{H}\circ\widetilde{\Psi}^{-1}$ on $\widetilde{\Omega}_{M}^{+}$ and the following diagram commutes.
$\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$$\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$$U^{+}$$U^{+}$$\scriptstyle{\tilde{H}}$$\scriptstyle{\widehat{\Pi}}$$\scriptstyle{\widehat{\Pi}}$$\scriptstyle{H}$
The next steps are repetitions of the techniques used in [15] and [6]. However, we briefly revisit them, since $H$ is a generalised Hénon map.
Proposition 2.11.
For every $\left[\frac{k}{d^{n}}\right]\in\uppi_{1}(U^{+})/\mathbb{Z}$, the deck transformation $\upgamma_{k/d^{n}}$ on $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ satisfies the relation
(2.11) $\widetilde{H}\circ\upgamma_{k/d^{n}}=\upgamma_{k/d^{n-1}}\circ\widetilde{H}$.
Thus, (2.2) holds.
Proof.
[…]
$\widehat{\upgamma}_{k/d^{n}}\left([p,C]\right)=\left[p,C_{k/d^{n}}\right]\text{ such that }\alpha\left(CC_{k/d^{n}}^{-1}\right)\in\left[\frac{k}{d^{n}}\right],$
where $\alpha$ is as defined in subsection 2.1. Let $\upgamma_{k/d^{n}}=\left(\widetilde{\Psi}\circ\widehat{\pi}\right)\circ\widehat{\upgamma}_{k/d^{n}}\circ\left(\widetilde{\Psi}\circ\widehat{\pi}\right)^{-1}.$ Fix a point $(z,\zeta)\in\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$, and let $\left(\widetilde{\Psi}\circ\widehat{\pi}\right)^{-1}(z,\zeta)=[p,C]$. Then
$\widetilde{H}\circ\upgamma_{k/d^{n}}(z,\zeta)$ $=\widetilde{H}\circ\widetilde{\Psi}\circ\widehat{\pi}\circ\widehat{\upgamma}_{k/d^{n}}\left([p,C]\right)=\widetilde{H}\circ\widetilde{\Psi}\circ\widehat{\pi}\left(\left[p,C_{k/d^{n}}\right]\right)$
$=\widetilde{\Psi}\circ\widehat{\pi}\left(\left[H(p),lH\left(C_{k/d^{n}}\right)\right]\right).$
Now
$\alpha\left(H\left(CC_{k/d^{n}}^{-1}\right)\right)=\alpha\left(lH\left(C\right)H\left(C_{k/d^{n}}^{-1}\right)l^{-1}\right)\in\left[\frac{k}{d^{n-1}}\right].$
Hence,
$\widetilde{H}\circ\upgamma_{k/d^{n}}(z,\zeta)=\widetilde{\Psi}\circ\widehat{\pi}\circ\widehat{\upgamma}_{k/d^{n-1}}\left(\left[H(p),lH\left(C\right)\right]\right)=\upgamma_{k/d^{n-1}}\circ\widetilde{H}(z,\zeta).$
For $n=1$, by comparing the coefficients in the above equation, it follows that
$\upgamma_{k/d}(z,\zeta)=\left(z+\frac{d}{a}\left(Q(\zeta)-Q(e^{2\pi ik/d}\zeta)\right),e^{2\pi ik/d}\zeta\right).$
Thus using (2.11) inductively (as in [6, page 24-25]), the deck transformations are given by
$\upgamma_{k/d^{n}}(z,\zeta)=\left(z+\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right),e^{2\pi ik/d^{n}}\cdot\zeta\right).\qed$
3. ${\rm Aut}(U^{+})$ and ${\rm Aut}(\Omega_{c}^{\prime})$
The goal of this section is to understand the structure of ${\rm Aut}_{1}(U^{+})$. Note that maps in this subgroup admit lifts to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ and it is precisely this property that makes them amenable for further classification. Furthermore, we will use this observation, to compute the automorphism group of the punctured Short $\mathbb{C}^{2}$’s, arising as the sub-level set of the Green’s function of $H$, consisting of maps that induce identity on the fundamental group. Finally, patching these observations leads to a description of the automorphism groups of both $U^{+}$ and punctured Short $\mathbb{C}^{2}$’s.
3.1. The subgroup ${\rm Aut}_{1}(U^{+})$
As mentioned before, we adapt the idea of Bousch [8] to prove the theorem given below. Similar ideas were also used to describe the automorphism group of punctured Short $\mathbb{C}^{2}$’s in [6] for simple Hénon maps. Recall the polynomial $Q$, as constructed in the proof of Theorem 2.1 (also see Proposition 2.7).
Theorem 3.1.
Let ${\rm Aut}_{1}(U^{+})$ be as above. Then
$\mathbb{C}\leq{\rm Aut}_{1}(U^{+})\leq\mathbb{C}\rtimes\mathbb{Z}_{d_{0}(d-1)},$
where $d_{0}$ is an appropriate factor of $d+d^{\prime}$.
Proof.
The proof consists of several steps. The main ingredient is to understand the lifts of elements in ${\rm Aut}_{1}(U^{+})$ to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ in a progressively better manner.
Step 1: Let $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$. First, we claim that any lift of the map $\mathfrak{a}$ to the covering space $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ is of the form
$A(z,\zeta)=(\beta z+\gamma(\zeta),\alpha\zeta),$
where $\alpha,\beta\in\mathbb{C}$ with $|\alpha|,|\beta|=1$ and $\gamma$ is holomorphic. Furthermore, for every pair of integers $k,n$, the lift $A$ commutes with the deck transformation $\upgamma_{k/d^{n}}$, i.e.,
$A\circ\upgamma_{k/d^{n}}(z,\zeta)=\upgamma_{k/d^{n}}\circ A(z,\zeta).$
To see this, let $A(z,\zeta)=\left({\rm A}_{1}(z,\zeta),{\rm A}_{2}(z,\zeta)\right).$ For a fixed $\zeta\in\mathbb{C}\setminus\overline{\mathbb{D}}$, the map ${\rm A}_{2}(\cdot,\zeta)$ is an entire holomorphic function whose range is contained in $\mathbb{C}\setminus\overline{\mathbb{D}}$. Hence, ${\rm A}_{2}(\cdot,\zeta)$ is a constant map and we can write ${\rm A}_{2}(z,\zeta)=h(\zeta)$. As $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$, both $\mathfrak{a}$ and $\mathfrak{a}^{-1}$ induce the identity on the fundamental group of $U^{+}$ and hence admit lifts to the covering space $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. Let $A^{\prime}$ be a lift of $\mathfrak{a}^{-1}$. Then
$\widehat{\Pi}\circ A\circ A^{\prime}(z,\zeta)=\mathfrak{a}\circ\widehat{\Pi}\circ A^{\prime}(z,\zeta)=\mathfrak{a}\circ\mathfrak{a}^{-1}\circ\widehat{\Pi}(z,\zeta)=\widehat{\Pi}(z,\zeta),$
where $\widehat{\Pi}$ is the covering map from $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ to $U^{+}$ as in Subsection 2.4. Note that, both $A\circ A^{\prime}$ and $A^{\prime}\circ A$ are lifts of the identity map. Hence, there exist integers $k,k^{\prime},n,n^{\prime}$ such that
[…]
Since $\upgamma_{k/d^{n}}$ and $\upgamma_{k^{\prime}/d^{n^{\prime}}}$ are deck transformations of the form (2.2), both $A$ and $A^{\prime}$ are automorphisms of $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. Let
$A^{-1}=\left(\tilde{{\rm A}}_{1}(z,\zeta),\tilde{h}(\zeta)\right).$
Thus, $\tilde{{\rm A}}_{1}\circ(A_{1}(\cdot,\zeta),h(\zeta))=\textsf{Id}$ for every $\zeta\in\mathbb{C}\setminus\overline{\mathbb{D}}$ and $h\circ\tilde{h}\equiv\tilde{h}\circ h\equiv\textsf{Id}$. Hence, ${\rm A}_{1}(\cdot,\zeta)$ is an automorphism of $\mathbb{C}$ and $h$ is an automorphism of $\mathbb{C}\setminus\overline{\mathbb{D}}$. In particular, there exists an $\alpha$ with $|\alpha|=1$ such that
(3.1) $h(\zeta)=\alpha\zeta\text{ and }{\rm A}_{1}(z,\zeta)=\beta(\zeta)z+\gamma(\zeta).$
Now, choose $(z^{\prime},\zeta^{\prime})$ and $(z,\zeta)$ in the same fibre. Since $A$ is a lift of the automorphism $\mathfrak{a}$, $A(z,\zeta)$ and $A(z^{\prime},\zeta^{\prime})$ also lie on the same fibre. Thus, by Theorem 2.1, there exist deck transformations $\upgamma_{k/d^{n}}$ and $\upgamma_{k^{\prime}/d^{n^{\prime}}}$ such that
$\upgamma_{k/d^{n}}\circ A(z,\zeta)=A(z^{\prime},\zeta^{\prime})=A\circ\upgamma_{k^{\prime}/d^{n^{\prime}}}(z,\zeta).$
By comparing the second coordinate in the above equation, it follows that $\left[\frac{k}{d^{n}}\right]=\left[\frac{k^{\prime}}{d^{n^{\prime}}}\right]$ and hence
$\upgamma_{k/d^{n}}\circ A(z,\zeta)=A\circ\upgamma_{k/d^{n}}(z,\zeta).$
Also, from (2.2),
$\upgamma_{k/d^{n}}(z,\zeta)=\left(z+\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right),e^{2\pi ik/d^{n}}\cdot\zeta\right)=(z^{\prime},\zeta^{\prime}).$
Hence $\zeta^{\prime}=e^{2\pi ik/d^{n}}\zeta$, and by equating the first coordinate, we get
$z+\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right)=z^{\prime}.$
Again, by equating the first coordinate of $\upgamma_{k/d^{n}}\circ A(z,\zeta)=A(z^{\prime},\zeta^{\prime})$, we get
$\beta(\zeta)z+\beta(\zeta)\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\alpha^{d^{l}}\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\alpha\zeta\right)^{d^{l}}\right)\right)+\gamma(\zeta)=\beta(\zeta^{\prime})z^{\prime}+\gamma(\zeta^{\prime}).$
Note that the above identities are true whenever both $(z,\zeta)$ and $(z^{\prime},\zeta^{\prime})$ lie in the same fibre. Thus, by comparing coefficients of $z$ on both sides, $\beta(\zeta)=\beta(\zeta^{\prime})=\beta(e^{2\pi ik/d^{n}}\zeta)$, for all integers $k,n\geq 1$. That is, $\beta$ is constant on the circle of radius $|\zeta|>1$. Hence, $\beta$ is a constant function. So, the above relation reduces to the following
$\beta\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right)+\gamma(\zeta^{\prime})$
$=\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\alpha^{d^{l}}\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\alpha\zeta\right)^{d^{l}}\right)\right)+\gamma(\zeta).$
Also, $\zeta/\zeta^{\prime}$ lies on the unit circle whenever $(z^{\prime},\zeta^{\prime})$ is on the same fibre as $(z,\zeta)$. So, $\gamma(\zeta)-\gamma(\zeta^{\prime})$ is uniformly bounded for a fixed $\zeta$. Thus, for every $k,n\geq 1$,
$\beta\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right)-\frac{d}{a}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l}\left(Q\left(\alpha^{d^{l}}\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\alpha\zeta\right)^{d^{l}}\right)\right)$
(3.2) $=$ $\left(\frac{d}{a}\right)^{n}\left[\left(\beta-\alpha^{d^{n-1}(d+d^{\prime})}\right)\left(1-e^{2\pi ikd^{\prime}/d}\right)\zeta^{d^{n-1}(d+d^{\prime})}\right]+R_{k,n}(\zeta)=\gamma(e^{2\pi ik/d^{n}}\zeta)-\gamma(\zeta),$
where $R_{k,n}(\zeta)$ is a polynomial of degree at most $d^{n}$. Let $a_{n}=\alpha^{d^{n-1}(d+d^{\prime})}.$
Claim: $a_{n}=\beta$ for $n$ sufficiently large.
The proof of the claim is the same as discussed in [6, Page 27]. We briefly revisit the steps. Fix a $k=1$. Let
(3.3) $L_{n}(\zeta)=Q(\zeta)-Q(e^{2\pi i/d^{n}}\zeta)\text{ and }\widetilde{L}_{n}^{l}(\zeta)=\beta L_{n-l}(\zeta)-L_{n-l}(\alpha^{d^{l}}\zeta)$
where $n\geq 1$ and $0\leq l\leq n-1$. Then (3.1) can be expressed as
(3.4) $\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l+1}\widetilde{L}_{n}^{l}\left(\zeta^{d^{l}}\right)=\gamma\left(e^{2\pi i/d^{n}}\zeta\right)-\gamma(\zeta).$
Since $Q$ is a monic polynomial of degree $d+d^{\prime}$ and $|\alpha|=1$, both $L_{n}$ and $\widetilde{L}_{n}^{l}$ are polynomials of degree at most $d+d^{\prime}$ and their coefficients are uniformly bounded for every $n\geq 1$ and $0\leq l\leq n-1$. Let
$\widetilde{L}_{n}^{l}(\zeta)=M_{n}^{l}\zeta^{d+d^{\prime}}+\text{ l.o.t.}$
Then choose $M>1$ such that
$\left|\widetilde{L}_{n}^{l}(\zeta)\right|<M|\zeta|^{d+d^{\prime}}\text{ and }\left|\widetilde{L}_{n}^{l}(\zeta)-M_{n}^{l}\zeta^{d+d^{\prime}}\right|<M|\zeta|^{d+d^{\prime}-1}$
for $|\zeta|>1$. Note that $M_{n}^{n-1}=(\beta-a_{n})(1-e^{2\pi id^{\prime}/d})$ and
$a_{n}=\left(\alpha^{d+d^{\prime}}\right)^{d^{n-1}}=f^{n-1}\left(\alpha^{d+d^{\prime}}\right)$
where $f(z)=z^{d}$. Thus, either $\{a_{n}\}$ is dense in the unit circle or $a_{n}=a_{n}^{d^{l}}=a_{n+l}$ for some $n\geq n_{0}$ and $l\geq 1$, i.e., $a_{0}$ is a pre-periodic fixed point of $f$. Suppose that the sequence $\{a_{n}\}$ does not converge to $\beta$. Then in either of the above cases, there exists a convergent subsequence say $\{a_{n_{l}}\}$ such that for every $l\geq 1$,
$|a_{n_{l}}-\beta|>\widetilde{\delta},\text{ i.e., }\lim_{l\to\infty}M_{n_{l}}^{n_{l}-1}=\delta\neq 0.$
Now from (3.3) and (3.4),
$\left|\gamma\left(e^{2\pi i/d^{n}}\zeta\right)-\gamma(\zeta)-\left(\frac{d}{a}\right)^{n}M_{n}^{n-1}\zeta^{d^{n-1}(d+d^{\prime})}\right|\leq\max\left\{1,\left|\frac{d}{a}\right|^{n}\right\}nM|\zeta|^{d^{n-1}(d+d^{\prime}-1)}.$
for $|\zeta|>1$. By the ratio test, for a fixed $|\zeta|>1$, both $\zeta^{d^{n-1}}\to\infty$ and $\left(\frac{d}{a}\right)^{n}\zeta^{d^{n-1}}\to\infty$ as $n\to\infty.$ Also $\gamma\left(e^{2\pi i/d^{n}}\zeta\right)-\gamma(\zeta)$ is bounded for every $n\geq 1$. Hence, we have
$0\neq|\delta|=\lim_{l\to\infty}\left|M_{n_{l}}^{n_{l}-1}\right|=$ $\lim_{l\to\infty}\left|\frac{\gamma\left(e^{2\pi i/d^{n_{l}}}\zeta\right)-\gamma(\zeta)}{\left(\frac{d}{a}\right)^{n_{l}}\zeta^{d^{n_{l}-1}(d+d^{\prime})}}-M_{n_{l}}^{n_{l}-1}\right|$
$\leq$ $\lim_{l\to\infty}\max\left\{1,\left|\frac{d}{a}\right|^{n_{l}}\right\}\left|\frac{a}{d}\right|^{n_{l}}\frac{n_{l}M}{\left|\zeta^{d^{n_{l}-1}}\right|}=0,$
which is a contradiction! Thus $a_{n}\to\beta$, i.e., $a_{0}$ is pre-periodic. In particular, $\beta=\alpha^{d^{n}(d+d^{\prime})}=\alpha^{d^{n+1}(d+d^{\prime})}=\beta=\beta^{d}$ for $n\geq n_{0}$. Hence $\beta^{d-1}=1\text{ and }\alpha^{d^{n_{0}}(d+d^{\prime})(d-1)}=1.$
Let $d=\mathbf{p}_{1}^{m_{1}}\mathbf{p}_{2}^{m_{2}}\ldots\mathbf{p}_{l}^{m_{l}}$ for primes $\mathbf{p}_{i},1\leq i\leq l$ and $m_{i}\geq 1$. Note that
$\mathbb{Z}\left[\frac{1}{d}\right]=\left\{k\mathbf{p}_{1}^{n_{1}}\mathbf{p}_{2}^{n_{2}}\ldots\mathbf{p}_{l}^{n_{l}}:k\in\mathbb{Z},n_{i}\in\mathbb{Z},1\leq i\leq l\right\}.$
Define
(3.5) $d_{0}=\inf\left\{(d+d^{\prime})r\in\mathbb{N}:r=\mathbf{p}_{1}^{-n_{1}}\mathbf{p}_{2}^{-n_{2}}\ldots\mathbf{p}_{l}^{-n_{l}},n_{i}\in\mathbb{N}\cup\{0\},1\leq i\leq l\right\}.$
Note that $(d+d^{\prime})d_{0}^{-1}\in\mathbb{Z}\left[\frac{1}{d}\right]$ and is a positive integer. In particular, $d_{0}$ is a factor of $(d+d^{\prime})$. Observe that if $H$ is a simple Hénon map then $d^{\prime}=1$ and $d_{0}=d+1$.
Step 2: For every $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$ there exists a unique lift to the covering space $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ of the form
$\widetilde{A}(z,\zeta)=(\tilde{\beta}z+\tilde{\gamma}(\zeta),\tilde{\alpha}\zeta),$
where $\tilde{\alpha}^{d^{n-1}(d+d^{\prime})}=\tilde{\beta}$ for every $n\geq 1$ and $\tilde{\beta}^{d-1}=\tilde{\alpha}^{d_{0}\big(d-1\big)}=1$.
To see this, let $A$ be an arbitrary lift of the automorphism $\mathfrak{a}$. By Step 1,
$A(z,\zeta)=(\beta z+\gamma(\zeta),\alpha\zeta).$
Then by the above step there exists $k_{0}\geq 1$ and $n_{0}\geq 1$ such that
$\alpha=e^{{2\pi ik_{0}}/{d^{n_{0}-1}}(d+d^{\prime})(d-1)}.$
Since none of the $\mathbf{p}_{i},1\leq i\leq l$ can be the factors of both $d_{0}$ and d1d-1, it follows that $d_{0}(d-1)$ is coprime to $(d+d^{\prime})d_{0}^{-1}$ and $d^{n}$ for every $n\geq 1$. Thus, there exist integers $p,q$ such that
$pd^{n_{0}-1}(d+d^{\prime})d_{0}^{-1}+q(d-1)d_{0}=k_{0}.$
In particular,
$\frac{k_{0}}{d^{n_{0}-1}(d+d^{\prime})(d-1)}=\frac{q}{d^{n_{0}-1}(d+d^{\prime})d_{0}^{-1}}+\frac{p}{d_{0}(d-1)}.$
Now by construction of $d_{0}$,
$\frac{q^{\prime}}{d^{n_{0}^{\prime}}}=\frac{q}{d^{n_{0}-1}(d+d^{\prime})d_{0}^{-1}}\in\mathbb{Z}\left[\frac{1}{d}\right],$
i.e., there exists a deck transformation $\upgamma_{q^{\prime}/d^{n_{0}^{\prime}}}$. Thus, $\widetilde{A}=A\circ\upgamma_{q^{\prime}/d^{n_{0}^{\prime}}}^{-1}$ is a lift of the automorphism $\mathfrak{a}$, which satisfies the properties stated in Step 2.
To prove the uniqueness of the lift, let $\widehat{A}$ and $\widetilde{A}$ be lifts of $\mathfrak{a}$ such that
[…]
with $\hat{\alpha}^{d_{0}(d-1)}={\tilde{\alpha}}^{d_{0}(d-1)}=1=\hat{\beta}^{d-1}=\tilde{\beta}^{d-1}$, and $\hat{\alpha}^{(d+d^{\prime})}=\hat{\beta}$, $\tilde{\alpha}^{(d+d^{\prime})}=\tilde{\beta}$. Since $\widehat{A}\circ\widetilde{A}^{-1}$ is a lift of the identity map, it corresponds to a deck transformation. Assume,
$\alpha=e^{2\pi ik_{1}/d_{0}(d-1)}\text{ and }\tilde{\alpha}=e^{2\pi ik_{2}/d_{0}(d-1)}\zeta$
for some integers $k_{1},k_{2}$.Thus
$\widehat{A}\circ\widetilde{A}^{-1}(z,\zeta)=\left(\hat{\beta}\tilde{\beta}^{-1}z+h(\zeta),e^{2\pi i(k_{1}-k_{2})/d_{0}(d-1)}\zeta\right),$
i.e., $\beta=\tilde{\beta}$ and $\frac{k_{1}-k_{2}}{d_{0}(d-1)}\in\mathbb{Z}\left[\frac{1}{d}\right]$. Since $d_{0}(d-1)$ is coprime to $d$, it follows that $k_{1}-k_{2}\in\mathbb{Z}$ and consequently, $\widehat{A}\circ\widetilde{A}^{-1}=\textsf{Identity}$. This completes Step 2.
Step 3: Let $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$ be such that it admits a lift of the form
$A(z,\zeta)=(z+\gamma(\zeta),\zeta).$
Then $\gamma\equiv c$ for some $c\in\mathbb{C}$, i.e., $\gamma$ is a constant function.
To see this, note that since $A\circ\upgamma_{k/d^{n}}=\upgamma_{k/d^{n}}\circ A$ for every integer $k,n$, it follows that for a fixed $\zeta\in\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$,
$\gamma(e^{2\pi ik/d^{n}}\zeta)=\gamma(\zeta).$
Thus, $\gamma$ is constant on the circle of radius $|\zeta|$, and is therefore identically constant on $\mathbb{C}\setminus\overline{\mathbb{D}}$.
Step 4: Let $A=(z+c,\zeta)$, $c\in\mathbb{C}$. Then there exists $\mathfrak{a}\in\textrm{Aut}_{1}(U^{+})$ such that it admits $A$ as a lift.
Define the map $\mathfrak{a}$ on $U^{+}$ as
$\mathfrak{a}\circ\widehat{\Pi}(z,\zeta)=\widehat{\Pi}\circ A(z,\zeta).$
where $\widehat{\Pi}$ is the covering map from $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ to $U^{+}.$ Note that $\mathfrak{a}$ is well defined if the image of points in the same fibre is the same. To check this, let $(z,\zeta)$ and $(z^{\prime},\zeta^{\prime})$ be points in the same fibre. Then there exist $k,n\in\mathbb{Z}$ such that $\upgamma_{k/d^{n}}(z,\zeta)=(z^{\prime},\zeta^{\prime})$. As $A\circ\upgamma_{k/d^{n}}=\upgamma_{k/d^{n}}\circ A$
$\mathfrak{a}\circ\widehat{\Pi}(z^{\prime},\zeta^{\prime})=\widehat{\Pi}\circ A\circ\upgamma_{k/d^{n}}(z,\zeta)=\widehat{\Pi}\circ\upgamma_{k/d^{n}}\circ A(z,\zeta)=\mathfrak{a}\circ\widehat{\Pi}(z,\zeta).$
Thus, by Step 3 and Step 4, $\mathbb{C}\leq{\rm Aut}_{1}(U^{+})$.
Step 5: $\mathbb{C}\leq\textrm{Aut}_{1}(U^{+})\leq\mathbb{C}\rtimes\mathbb{Z}_{d_{0}(d-1)}$. In particular, there exists a subgroup $G$ of $\mathbb{Z}_{d_{0}(d-1)}$ and an operation
$\bullet:(\mathbb{C}\rtimes G)$ $\times(\mathbb{C}\rtimes G)\to\mathbb{C}\rtimes G$
$(\gamma_{1},\alpha_{1})$ $\bullet(\gamma_{2},\alpha_{2})=\left(\gamma_{1}+\phi_{\alpha_{1}}(\gamma_{2}),\alpha_{1}\alpha_{2}\right),$
where $\phi_{\alpha}:\mathbb{C}\to\mathbb{C}$ is a group isomorphism for every $\alpha\in G$ such that
$\left({\rm Aut}_{1}(U^{+}),\circ\right)\cong(\mathbb{C}\rtimes G,\bullet).$
To prove this, let $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$. By Step 2 there exists a unique lift
$A(z,\zeta)=(\beta z+\gamma(\zeta),\alpha\zeta),$
such that $\alpha^{d_{0}(d-1)}=\beta^{d-1}=1$. Fix a base point $\zeta_{0}$ in $\mathbb{C}\setminus\overline{\mathbb{D}}$ and define the map
$\mathfrak{G}:{\rm Aut}_{1}(U^{+})\to\mathbb{C}\times\mathbb{Z}_{d_{0}(d-1)}$
by $\mathfrak{G}(\mathfrak{a})=(\gamma(\zeta_{0}),\alpha)$ where $A$ is the lift as assumed above. Note that $\mathfrak{G}$ is well-defined, since by Step 2 the map $A$ is a unique lift of the automorphism $\mathfrak{a}$.
Suppose $\mathfrak{G}(\mathfrak{a})=\mathfrak{G}(\mathfrak{a}^{\prime})=(c,\alpha).$ Then there exist lifts of $\mathfrak{a}$ and $\mathfrak{a}^{\prime}$, say $A$ and $A^{\prime}$ such that
[…]
with $\gamma(\zeta_{0})=\gamma^{\prime}(\zeta_{0})=c$. Also, there exists a lift of the map $\mathfrak{a}^{\prime}\circ\mathfrak{a}^{-1}$ of the form
$A^{\prime}\circ A^{-1}(z,\zeta)=(z+\gamma^{\prime}(\alpha^{-1}\zeta)-\gamma(\alpha^{-1}\zeta),\zeta).$
By Step 3, $\gamma^{\prime}(\alpha^{-1}\zeta)-\gamma(\alpha^{-1}\zeta)\equiv c$ is a constant function. However, by assumption $\gamma(\zeta_{0})=\gamma^{\prime}(\zeta_{0})$, i.e., $A^{\prime}\circ A^{-1}(z,\alpha\zeta_{0})=(z,\alpha\zeta_{0})$. Hence $c=0$, or equivalently $\mathfrak{a}=\mathfrak{a}^{\prime}$.
Let
[…]
Furthermore, note that if $\alpha,\tilde{\alpha}\in G$ then both $\alpha^{-1}$ and $\alpha\tilde{\alpha}$ are in $G$, and hence $G$ is a subgroup of $\mathbb{Z}_{d_{0}(d-1)}$. Also, by Step 4, $\mathbb{C}\times G=\mathfrak{G}\left({\rm Aut}_{1}(U^{+})\right)$. Thus, $\mathfrak{G}$ induces the semi-direct product $\mathbb{C}\rtimes G$, and this completes the proof. ∎
3.2. The group ${\rm Aut}_{1}(\Omega_{c}^{\prime})$
Recall that
$\Omega_{c}=\{z\in\mathbb{C}^{2}:G_{H}^{+}(z)<c\},\;\;c>0$
is a Short $\mathbb{C}^{2}$. A punctured Short $\mathbb{C}^{2}$, denoted by $\Omega_{c}^{\prime}$, is the complement of the escaping set in $\Omega_{c}$, i.e., $\Omega_{c}^{\prime}=\Omega_{c}\setminus K^{+}$. Next, we give a generalised version of [6, Theorem 1.2] using the fact that the covering maps over punctured Short $\mathbb{C}^{2}$’s are restrictions of $\widehat{\Pi}$ on appropriate invariant subspaces of $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. As before, we will use ${\rm Aut}_{1}(\Omega_{c}^{\prime})$ to denote the automorphisms of $\Omega_{c}^{\prime}$ that induce the identity on the fundamental group of $\Omega_{c}^{\prime}$. The factor $d_{0}$ as in (3.5) will continue to appear in the sequel.
Theorem 3.2.
Let $\mathcal{A}_{c}=\{z\in\mathbb{C}:1<|z|<e^{c}\}$. The map $\widehat{\Pi}$ restricted to $\mathbb{C}\times\mathcal{A}_{c}$ for every $c>0$ is a covering over $\Omega_{c}^{\prime}$ and the following diagram commutes
$\mathbb{C}\times\mathcal{A}_{c}$$\mathbb{C}\times\mathcal{A}_{dc}$$\mathbb{C}\times\mathcal{A}_{d^{2}c}$$\cdots$$\Omega_{c}^{\prime}$$\Omega_{dc}^{\prime}$$\Omega_{d^{2}c}^{\prime}$$\cdots$$\scriptstyle{\tilde{H}}$$\scriptstyle{\widehat{\Pi}}$$\scriptstyle{\tilde{H}}$$\scriptstyle{\widehat{\Pi}}$$\scriptstyle{\tilde{H}}$$\scriptstyle{\widehat{\Pi}}$$\scriptstyle{H}$$\scriptstyle{H}$$\scriptstyle{H}$
Furthermore,
  1. (i)
    The fundamental group $\uppi_{1}(\Omega_{c}^{\prime})$ is isomorphic to $\mathbb{Z}\left[\frac{1}{d}\right]$ for every $c>0$.
  2. (ii)
    An element $\left[\frac{k}{d^{n}}\right]$ of $\uppi_{1}(\Omega^{\prime}_{c})/\mathbb{Z}$ corresponds to the following deck transformation of $\widehat{\Omega^{\prime}_{c}}$:
    $\upgamma_{k/d^{n}}(z,\zeta)=\left(z+\frac{d}{a}\sum_{l=0}^{\infty}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right),e^{2\pi ik/d^{n}}\cdot\zeta\right).$
  3. (iii)
    Also, $\mathbb{C}\leq{\rm Aut}_{1}(\Omega_{c}^{\prime})\leq\mathbb{C}\rtimes\mathbb{Z}_{d_{0}(d-1)}$.
Proof.
Recall that $\uppi_{1}(U^{+})=\mathbb{Z}\left[\frac{1}{d}\right]$. Let
$\mathcal{D}=\left\{\upgamma_{{k}/{d^{n}}}:\left[\frac{k}{d^{n}}\right]\in\mathbb{Z}\left[\frac{1}{d}\right]/\mathbb{Z}\right\}$
be the group of deck transformations. Note that $\upgamma_{k/d^{m}}(\mathbb{C}\times\mathcal{A}_{c})=\mathbb{C}\times\mathcal{A}_{c}.$ Also, the action of $\mathcal{D}$ is properly discontinuous on $\mathbb{C}\times\mathcal{A}_{c}.$ Hence
$(\mathbb{C}\times\mathcal{A}_{c})/\mathcal{D}\simeq\widehat{\Pi}(\mathbb{C}\times\mathcal{A}_{c}).$
Let $\tilde{\Omega}_{c}=\widehat{\Pi}(\mathbb{C}\times\mathcal{A}_{c})$. Thus $\mathbb{C}\times\mathcal{A}_{c}$ is a covering of $\tilde{\Omega}_{c}$ via the map $\widehat{\Pi}$ and $\uppi_{1}(\tilde{\Omega}_{c})=\mathbb{Z}\left[\frac{1}{d}\right]$. Next, we use Proposition 2.3 to prove that:
Claim: $\tilde{\Omega}_{c}=\Omega_{c}^{\prime}$ for every $c>0$.
Let $z\in\tilde{\Omega}_{c}$ and choose $p=[z,C]\in\widehat{U}^{+}$. Since $z\in U^{+}$, $H^{n}(z)\in V_{R}^{+}$ and $H^{n}(C)\subset V_{R}^{+}$, i.e., $\widehat{H}^{n}(p)\in\widehat{V}^{+}$ and $\tilde{H}^{n}(\mathbb{C}\times\mathcal{A}_{c})=\mathbb{C}\times\mathcal{A}_{cd^{n}}$ for every large enough $n\geq 1$. Hence, $\widehat{\Phi}(\widehat{H}^{n}(p))\in\mathbb{C}\times\mathcal{A}_{cd^{n}}.$ Thus
$\left|\hat{\phi}(\widehat{H}^{n}(p))\right|<e^{cd^{n}}\text{ or }\left|\hat{\phi}(p)^{d^{n}}\right|<e^{cd^{n}}.$
In particular, $|\hat{\phi}(p)|<e^{c}$. Also, $\hat{\phi}(\widehat{H}^{n}(p))=\phi(H^{n}(z))$ by Proposition 2.3 (i) and
$G_{H}^{+}(H^{n}(z))=\log|\phi(H^{n}(z))|<d^{n}c$
for every $n\geq 1$ large enough. Thus $G_{H}^{+}(z)<c$ for every $z\in\tilde{\Omega}_{c}$, and consequently we have $\tilde{\Omega}_{c}\subset\Omega_{c}^{\prime}$.
Conversely, let $z\in\Omega_{c}^{\prime}$ and choose $p=[z,C]\in\widehat{U}^{+}$. Note that for sufficiently large $n$,
$G_{H}^{+}(H^{n}(z))=\log|\phi(H^{n}(z))|<d^{n}c,$
or equivalently $G_{H}^{+}(z)<c$. Hence again by 2.3 (i), $\widehat{\Phi}(\widehat{H}^{n}(p))\in\mathbb{C}\times\mathcal{A}_{d^{n}c}$, or equivalently $\tilde{H}^{n}\circ\widehat{\Phi}(p)\in\mathbb{C}\times\mathcal{A}_{d^{n}c}$. Thus, $\widehat{\Phi}(p)\in\mathbb{C}\times\mathcal{A}_{c}$ and $z\in\widehat{\Pi}(\mathbb{C}\times\mathcal{A}_{c})=\tilde{\Omega}_{c}$.
Further, $\uppi_{1}(\Omega_{c}^{\prime})=\mathbb{Z}\left[\frac{1}{d}\right]$, and for every $\frac{k}{d^{n}}]\in\pi_{1}(\Omega_{c}^{\prime})/\mathbb{Z}$
$\widehat{\Pi}\circ\upgamma_{k/d^{n}}=\widehat{\Pi}$,
where $\upgamma_{k/d^{n}}$ is as in Theorem 2.1. Hence, the deck transformations of $\mathbb{C}\times\mathcal{A}_{c}$ are given by
$\upgamma_{k/d^{n}}(z,\zeta)=\left(z+\frac{d}{a}\sum_{l=0}^{\infty}\left(\frac{d}{a}\right)^{l}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right),e^{2\pi ik/d^{n}}\cdot\zeta\right).$
Now, by the reasoning used in Theorem 1.2 in [6] and Corollary 3.5 above, the statement on ${\rm Aut}_{1}(\Omega_{c}^{\prime})$ holds. It is indeed also similar to the proof of Theorem 3.1; however, we will summarize the steps for the sake of completeness.
Let $c>0$. By the reasoning in Step 1 in the proof of Theorem 3.1, the lift of any automorphism $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ to $\mathbb{C}\times\mathcal{A}_{c}$ has the form
$A(z,\zeta)=\left(\tilde{\beta}(\zeta)z+\tilde{\gamma}(\zeta),h(\zeta)\right)$
where $\tilde{\gamma},\tilde{\beta}$ are holomorphic functions on $\mathcal{A}_{c}$ and $h$ is an automorphism of $\mathcal{A}_{c}$. Further, $\tilde{\beta}(\zeta)\neq 0$ for every $\zeta\in\mathcal{A}_{c}$.
Case 1: Suppose $h(\zeta)=\tilde{\alpha}e^{c}/\zeta$ for some $|\tilde{\alpha}|=1$. Now, choose $(z^{\prime},\zeta^{\prime})$ and $(z,\zeta)$ on the same fibre. Then, by assumption $A(z,\zeta)$ and $A(z^{\prime},\zeta^{\prime})$ lies on the same fibre. In particular, for every deck transformation $\upgamma_{k/d^{n}}$ there exists a corresponding transformation $\upgamma_{k^{\prime}/d^{n^{\prime}}}$ such that $\upgamma_{k/d^{n}}\circ A(z,\zeta)=A(z^{\prime},\zeta^{\prime})=A\circ\upgamma_{k^{\prime}/d^{n^{\prime}}}(z,\zeta)$. By comparing the second coordinate, it follows that $\left[\frac{k}{d^{n}}\right]=-\left[\frac{k^{\prime}}{d^{n^{\prime}}}\right]$ or $k+k^{\prime}=md^{n}$ for some integer $m\geq 1$. Hence,
$\upgamma_{k/d^{n}}\circ A(z,\zeta)=A\circ\upgamma_{(md^{n}-k)/d^{n}}(z,\zeta)\text{ or }\upgamma_{k/d^{n}}\circ A\circ\upgamma_{k/d^{n}}(z,\zeta)=A(z,\zeta).$
Again, by comparing the first coordinate of the above expression first we have
$\tilde{\beta}\left(e^{2\pi ik^{\prime}/d^{n}}\zeta\right)z=\tilde{\beta}(\zeta)z.$
Since the above is true for every pair of positive integers $k^{\prime},n\geq 1$, $\tilde{\beta}$ is a non-zero constant map. Further,
$\tilde{\beta}\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l+1}\left(Q\left(\zeta^{d^{l}}\right)-Q\left(\left(e^{2\pi ik^{\prime}/d^{n}}\cdot\zeta\right)^{d^{l}}\right)\right)+\tilde{\gamma}\left(e^{2\pi ik^{\prime}/d^{n}}\zeta\right)$
(3.6) $=$ $\sum_{l=0}^{n-1}\left(\frac{d}{a}\right)^{l+1}\left(Q\left({\tilde{\alpha}e^{c}/\zeta}^{d^{l}}\right)-Q\left(\left(e^{2\pi ik^{\prime}/d^{n}}\cdot{\tilde{\alpha}e^{c}/\zeta}\right)^{d^{l}}\right)\right)+\gamma(\zeta).$
Now, by a similar computation as in Step 1 of the proof of Theorem 3.1, for a fixed $\zeta\in\mathbb{C}\setminus\overline{\mathbb{D}}$, the left side above in (3.2) diverges to infinity. However, the right side in (3.2) is bounded, which is a contradiction. Hence, $h(\zeta)\neq\tilde{\alpha}e^{c}/\zeta$ for some $|\tilde{\alpha}|=1$.
Case 2: Suppose $h(\zeta)=\tilde{\alpha}\zeta$ for some $|\tilde{\alpha}|=1$, then by the exactly same reasoning as in Step 1 and Step 2 of the proof of Theorem 3.1, it follows that there exists a unique lift of every $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ to $\mathbb{C}\times\mathcal{A}_{c}$ of the form
(3.7) $A(z,\zeta)=\left(\beta z+\gamma(\zeta),\alpha\zeta\right),$
where $\alpha^{d+d^{\prime}}=\beta$, $\beta^{d-1}=\alpha^{d_{0}(d-1)}=1$ and $\gamma$ is a holomorphic function on $\mathcal{A}_{c}$. Here $d_{0}$ is as defined by (3.5) in the proof of Theorem 3.1.
In particular, as in Step 5 of Theorem 3.1, for every $c>0$ there exists an injective map from
$\mathfrak{G}_{c}:{\rm Aut}_{1}(\Omega_{c}^{\prime})\to\mathbb{C}\times\mathbb{Z}_{d_{0}(d-1)},$
given by $\mathfrak{G}_{c}(\mathfrak{a})=(\gamma(\zeta_{c}),\alpha)$ where $\zeta_{c}$ is a fixed point in $\mathcal{A}_{c}$. Now by repeating Steps 3,4 and 5 from the proof of Theorem 3.1, the result follows. ∎
Remark 3.3.
Let $f\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ such that the $\mathfrak{G}_{c}(f)=(a_{c},\alpha)$. Then for every $0<b<c$ and $(z,\zeta)\in\mathbb{C}\times\mathcal{A}_{b}$,
$f\circ\widehat{\Pi}(z,\zeta)=\widehat{\Pi}(\alpha^{d+d^{\prime}}z+\gamma(\zeta),\alpha\zeta)\text{ where }\gamma(\zeta_{c})=a_{c}.$
Hence $f\in{\rm Aut}_{1}(\Omega_{b}^{\prime})$ and $\mathfrak{G}_{b}(f)=(a_{b},\alpha)$ where $\gamma(\zeta_{b})=a_{b}$. In particular, if $\alpha=1$ then $\gamma$ is a constant function, hence $a_{c}=a_{b}$. Thus for every $a\in\mathbb{C}$ and $0<b<c$
$\mathfrak{G}_{c}^{-1}(a,1)=\mathfrak{G}_{b}^{-1}(a,1)=(z+a,\zeta).$
Similarly, $\mathfrak{G}^{-1}(a,1)=\mathfrak{G}_{c}^{-1}(a,1)$ for every $c>0$ and $a\in\mathbb{C}$, where $\mathfrak{G}$ is as defined in Step 5 of Theorem 3.1.
3.3. The automorphism groups of $U^{+}$ and $\Omega_{c}^{\prime}$
As a consequence of Theorem 3.1 and 3.2, we prove the following results on the entire automorphism group of $U^{+}$ and $\Omega_{c}^{\prime}$ for every $c>0.$
Theorem 3.4.
Let $H$ be a map of the form (1.1). Then
  • (i)
    $\mathbb{C}\times\mathbb{Z}\leq{\rm Aut}(U^{+})\leq\mathbb{C}\rtimes\mathbb{Z}_{d_{0}(d-1)}\times\mathbb{Z}$ if $d=\mathbf{p}^{m}$ for a prime number $\mathbf{p}$ and a positive integer $m\geq 1$.
  • (ii)
    $\mathbb{C}\times\mathbb{Z}\leq{\rm Aut}(U^{+})\leq\left(\mathbb{C}\rtimes\mathbb{Z}_{d_{0}(d-1)}\right)\times\mathbb{Z}^{l}$ if $d=\mathbf{p}_{1}^{m_{1}}\ldots\mathbf{p}_{l}^{m_{l}}$ for primes $\mathbf{p}_{i},1\leq i\leq l$, and $m_{i}\geq 1$.
Proof.
Let $d=\mathbf{p}_{1}^{m_{1}}\ldots\mathbf{p}_{l}^{m_{l}}$ for primes $\mathbf{p}_{i},1\leq i\leq l$, and $m_{i}\geq 1$. Recall that, the fundamental group of $U^{+}$ is
$\uppi_{1}(U^{+})\simeq\mathbb{Z}\left[\frac{1}{d}\right]=\left\{\frac{m}{d^{k}}:m,k\in\mathbb{Z}\right\},$
which is a commutative ring with the usual binary operations. We first recall the following straightforward:
Claim: The units of the ring $\mathbb{Z}\left[\frac{1}{d}\right]$ are isomorphic to $\mathbb{Z}^{l}\times\mathbb{Z}_{2}$. In particular,
$U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right)=\left\{\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}:n_{i}\in\mathbb{Z},1\leq i\leq l\right\}.$
Let $u=\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}$ for $n_{i}\in\mathbb{Z},1\leq i\leq l$. Let $k,\tilde{k}\geq 1$ such that $q_{i}=km_{i}+n_{i}\geq 0$ and $\tilde{q}_{i}=\tilde{k}m_{i}-n_{i}\geq 0.$ Then
$u=\pm\frac{\mathbf{p}_{1}^{q_{1}}\ldots\mathbf{p}_{l}^{q_{l}}}{d^{k}}\text{ and }u^{-1}=\pm\frac{\mathbf{p}_{1}^{\tilde{q}_{1}}\ldots\mathbf{p}_{l}^{\tilde{q}_{l}}}{d^{\tilde{k}}}$,
i.e., $u\in U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right)$. Now suppose $u\in U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right)$. Then there exist $m,m^{\prime},k,k^{\prime}\in\mathbb{Z}$ such that $u=\frac{m}{d^{k}}$ and $u^{-1}=\frac{m^{\prime}}{d^{k^{\prime}}}$. Thus, $mm^{\prime}=d^{k+k^{\prime}}$ or equivalently $m=\pm\mathbf{p}_{1}^{q_{1}}\ldots\mathbf{p}_{l}^{q_{l}}$ for positive integers $q_{i}\geq 0$. Hence, $u\in\{\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}:n_{i}\in\mathbb{Z},1\leq i\leq l\}$.
Let $\mathfrak{i}$ be a group isomorphism of $\mathbb{Z}\left[\frac{1}{d}\right]$. Since $\mathfrak{i}\left(\frac{m}{d^{k}}\right)=\frac{m}{d^{k}}\mathfrak{i}(1)$ for every $m,k\in\mathbb{Z}$ and $\mathfrak{i}(1).\mathfrak{i}^{-1}(1)=1$, $\mathfrak{i}(1)=u_{\mathfrak{i}}\in U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right).$ Hence $\mathfrak{i}$ is an isomorphism, if and only if there exists a unit $u_{\mathfrak{i}}\in U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right)$ such that $\mathfrak{i}(x)=u_{\mathfrak{i}}x$ for every $x\in\mathbb{Z}\left[\frac{1}{d}\right]$. For $\phi\in{\rm Aut}(U^{+})$, let $\mathfrak{i}_{\phi}(x)=u_{\mathfrak{i}_{\phi}}x$ denote the isomorphism induced by $\phi$ on $\mathbb{Z}\left[\frac{1}{d}\right]$. Let
$\tilde{I}=\left\{(n_{1},\ldots,n_{l})\in\mathbb{Z}^{l}:\text{ there exists }\phi\in{\rm Aut}(U^{+})\text{ such that }u_{\mathfrak{i}_{\phi}}=\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}\right\}.$
Thus $\tilde{I}\leq\mathbb{Z}^{l}\leq U\left(\mathbb{Z}\left[\frac{1}{d}\right]\right)$. Also $H\in{\rm Aut}(U^{+})$, and $\mathfrak{i}_{H}(x)=dx$ for every $x\in\mathbb{Z}\left[\frac{1}{d}\right]$. Hence, $\mathbb{Z}\simeq\{(m_{1}k,\ldots,m_{l}k):k\in\mathbb{Z}\}\leq\tilde{I}$. Now, consider the map $h:{\rm Aut}(U^{+})\to\tilde{I}$ such that $h(\phi)=(n_{1},\ldots,n_{l})$ where $u_{\mathfrak{i}_{\phi}}=\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}$, is a surjective group homomorphism with kernel $I_{0}=\{\phi\in{\rm Aut}(U^{+}):u_{\mathfrak{i}_{\phi}}=\pm 1\}.$ Suppose there exists $\phi\in I_{0}$ such that $u_{\mathfrak{i}_{\phi}}=-1$, then $\mathfrak{i}_{\phi}(\mathbb{Z})=\mathbb{Z}.$ In particular, $\phi$ lifts to an automorphism $\tilde{\phi}$ of the covering space $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. By the analysis in Step 1 of the proof of Theorem 3.1,
$\tilde{\phi}(z,\zeta)=(\beta z+\gamma(\zeta),\alpha\zeta)\text{ and }\mathfrak{i}_{\tilde{\phi}}(n)=n\text{ for every }n\in\mathbb{Z}\simeq\pi_{1}\left(\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)\right).$
Let $\widehat{\Pi}$ be the covering map from $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ to $U^{+}$, then $\widehat{\Pi}_{*}(n)=n$ for every $n\in\mathbb{Z}\simeq\uppi_{1}\left(\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)\right)$. But $\widehat{\Pi}_{*}\circ\mathfrak{i}_{\tilde{\phi}}(n)=\mathfrak{i}_{\phi}\circ\widehat{\Pi}_{*}(n)$, i.e., n=nn=-n for every $n\in\mathbb{Z}$. This is not possible, hence $u_{\mathfrak{i}_{\phi}}=1$ for every $\phi\in I_{0}$ or $I_{0}={\rm Aut}_{1}(U^{+})$ and ${\rm Aut}_{1}(U^{+})\times\tilde{I}\simeq{\rm Aut}(U^{+})$. Thus,
$\mathbb{C}\times\mathbb{Z}\leq{\rm Aut}_{1}(U^{+})\times\mathbb{Z}\leq{\rm Aut}_{1}(U^{+})\times\tilde{I}\simeq{\rm Aut}(U^{+})\leq{\rm Aut}_{1}(U^{+})\times\mathbb{Z}^{l}\leq\left(\mathbb{C}\rtimes\mathbb{Z}_{(d+d^{\prime})(d-1)}\right)\times\mathbb{Z}^{l}.$
This proves (ii). Now, if $d=\mathbf{p}^{m}$ for a prime $\mathbf{p}$ and positive integer $m\geq 1$, the above identity simplifies as
$\mathbb{C}\times\mathbb{Z}\leq{\rm Aut}(U^{+})\leq\left(\mathbb{C}\rtimes\mathbb{Z}_{(d+d^{\prime})(d-1)}\right)\times\mathbb{Z}.\qed$
Corollary 3.5.
$\mathbb{C}\leq{\rm Aut}(\Omega_{c}^{\prime})\leq\left(\mathbb{C}\rtimes\mathbb{Z}_{(d+d^{\prime})(d-1)}\right)\times\mathbb{Z}^{l-1}$ where $d=\mathbf{p}_{1}^{m_{1}}\ldots\mathbf{p}_{l}^{m_{l}}$ for primes $\mathbf{p}_{i},1\leq i\leq l$ and $m_{i}\geq 1$.
Proof.
By the proof of Theorem 3.2 (iii), for every $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ there exists a unique lift of the form $(\beta z+\gamma(\zeta),\alpha\zeta)$ such that $\alpha^{d+d^{\prime}}=\beta$ and $\beta^{d-1}=\alpha^{d_{0}(d-1)}=1$, where $d_{0}$ is as in (3.5). Also $\uppi_{1}(\Omega_{c}^{\prime})=\mathbb{Z}\left[\frac{1}{d}\right]$ and by repeating the same reasoning as in the proof of Theorem 3.4,
${\rm Aut}(\Omega_{c}^{\prime})\simeq{\rm Aut}_{1}(\Omega_{c}^{\prime})\times\tilde{I}_{c},$
where $\tilde{I}_{c}=\left\{(n_{1},\ldots,n_{l})\in\mathbb{Z}^{l}:\text{ there exists }\phi\in{\rm Aut}(\Omega_{c}^{\prime})\text{ such that }u_{\mathfrak{i}_{\phi}}=\pm\mathbf{p}_{1}^{n_{1}}\ldots\mathbf{p}_{l}^{n_{l}}\right\}.$
Claim: $\tilde{I}_{c}$ is isomorphic to a subgroup $\mathbb{Z}^{l-1}$ for every $c>0$.
Suppose not, i.e., $\tilde{I}_{c}\simeq\mathbb{Z}^{l}$. Then there are integers $r_{i}\geq 1$ such that
$\tilde{I}_{c}=r_{1}\mathbb{Z}\times\cdots\times r_{l}\mathbb{Z}.$
Let $r=r_{1}r_{2}...r_{l}$, then $(rm_{1},\ldots,rm_{k})\in\tilde{I}_{c}$, which means that there exists $\psi\in{\rm Aut}(\Omega_{c}^{\prime})$ such that $u_{\mathfrak{i}_{\psi}}=d^{r}$. Define $\mathfrak{a}=H^{-r}\circ\psi$. Then $\mathfrak{a}(\Omega_{c}^{\prime})=\Omega^{\prime}_{cd^{-r}}$, and $\mathfrak{a}$ is a biholomorphism that induces the identity map between $\uppi_{1}(\Omega_{c}^{\prime})$ and $\uppi_{1}\left(\Omega_{cd^{-r}}^{\prime}\right)$. Hence, $\mathfrak{a}$ lifts to a biholomorphism between $\mathbb{C}\times\mathcal{A}_{c}$ to $\mathbb{C}\times\mathcal{A}_{cd^{-r}}$. Let
$A(z,\zeta)=\left({\rm A}_{1}(z,\zeta),{\rm A}_{2}(z,\zeta)\right)$
be this biholomorphic lift. Now $z\mapsto{\rm A}_{2}(z,\zeta)$ is a bounded entire function, which implies that ${\rm A}_{2}(z,\zeta)\equiv h(\zeta)$ and thus $A(z,\zeta)=({\rm A}_{1}(z,\zeta),h(\zeta))$. Similarly, $A^{-1}(z,\zeta)=({\rm A}_{1}^{*}(z,\zeta),h^{*}(\zeta)).$ Since $A\circ A^{-1}(z,\zeta)=A^{-1}\circ A(z,\zeta))=(z,\zeta),$ it follows that $h\circ h^{*}\equiv h^{*}\circ h\equiv\sf{Id}$. Therefore, $h$ is a biholomorphism between $\mathcal{A}_{c}$ and $\mathcal{A}_{cd^{-r}}$, which is not possible, unless $r=0$. Hence, the claim follows.
Thus, $\mathbb{C}\leq{\rm Aut}_{1}(\Omega_{c}^{\prime})\leq{\rm Aut}(\Omega_{c}^{\prime})\simeq{\rm Aut}_{1}(\Omega_{c}^{\prime})\times\tilde{I}_{c}\leq(\mathbb{C}\rtimes\mathbb{Z}_{(d+d^{\prime})(d-1)})\times\mathbb{Z}^{l-1}.$
4. The subgroup ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ and Proof of Theorem 1.1
In this section, we will use the lifts of elements in ${\rm Aut}_{1}(U^{+})$ to describe the elements of ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$. The map $\mathfrak{G}$, introduced in the proof of Theorem 3.1, will continue to denote the isomorphism between ${\rm Aut}_{1}(U^{+})$ and $\mathbb{C}\rtimes G$. The term $\mathfrak{G}$-lift will be used to denote the lift of an automorphism of $U^{+}$ corresponding to the map $\mathfrak{G}$, i.e., for $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$ if $\mathfrak{G}(\mathfrak{a})=(c,\alpha)$ then the $\mathfrak{G}$-lift of $\mathfrak{a}$ is
$A(z,\zeta)=(\alpha^{d+d^{\prime}}z+\gamma(\zeta),\alpha\zeta)\text{ where }\gamma(\zeta_{0})=c.$
Furthermore, by Step 3 of the proof of Theorem 3.1, if $\mathfrak{G}(\mathfrak{a})=(c,1)$ then the lift of $\mathfrak{a}$ is $(z+c,\zeta)$. In particular, the lift is independent of the point $\zeta_{0}$ in $\mathbb{C}\rtimes\{1\}$.
4.1. The group $\mathfrak{G}^{-1}\left(\mathbb{C}\rtimes\{1\}\right)\cap{\rm Aut}(\mathbb{C}^{2})$
The goal of this subsection is to prove Theorem 4.1 which says that $\mathfrak{G}^{-1}\left(\mathbb{C}\rtimes\{1\}\right)\cap{\rm Aut}(\mathbb{C}^{2})=\textsf{Identity}$. The proof will be presented through various cases.
Theorem 4.1.
Let $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ be such that the $\mathfrak{G}$-lift of $\mathfrak{a}$ is of the form
$A(z,\zeta)=(z+c_{0},\zeta).$
Then $c_{0}=0$, or equivalently $\mathfrak{a}$ is the identity map.
Define the group isomorphisms $\Phi_{H}^{\pm}:{\rm Aut}(U^{+})\to{\rm Aut}(U^{+})$
(4.1) $\Phi_{H}^{+}(f)$ $=H\circ f\circ H^{-1}\text{ and }\Phi_{H}^{-}(f)=H^{-1}\circ f\circ H\text{ for }f\in{\rm Aut}(U^{+})$
and note that from Proposition 2.2(i), ${\rm Aut}_{1}(U^{+})$ is invariant under $\Phi_{H}^{\pm}$ with $\Phi_{H}^{+}\circ\Phi_{H}^{-}=\Phi_{H}^{-}\circ\Phi_{H}^{+}=\textsf{Identity}$. Recall from Theorem 2.1 that the lift of the map $H$ to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ is
$\widetilde{H}(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta),\zeta^{d}\right),$
where $Q$ is a polynomial of degree $d+d^{\prime}$.
Lemma 4.2.
The $\mathfrak{G}$-lifts of the automorphisms $\Phi_{H}^{\pm}(\mathfrak{a})$, denoted by $\widetilde{\Phi_{H}^{\pm}(\mathfrak{a})}$, are of the form
[…]
where $\mathfrak{a}\in\mathfrak{G}^{-1}\left(\mathbb{C}\rtimes\{1\}\right)$.
Proof.
Note that $H\circ\mathfrak{a}=\Phi_{H}^{+}(\mathfrak{a})\circ H$. Hence, there exists a lift of $H\circ\mathfrak{a}$ to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ of the form
$\widetilde{H\circ\mathfrak{a}}(z,\zeta)=\widetilde{H}\circ A(z,\zeta)=\left(\frac{a}{d}z+Q(\zeta)+\frac{a}{d}c_{0},\zeta^{d}\right)=A^{\prime}\circ\widetilde{H}(z,\zeta),$
where $A^{\prime}(z,\zeta)=\left(z+\frac{a}{d}c_{0},\zeta\right)$. Now by Theorem 3.4, there exists $\mathfrak{a}^{\prime}\in{\rm Aut}_{1}(U^{+})$ such that $A^{\prime}$ corresponds to the $\mathfrak{G}$-lift of $\mathfrak{a}^{\prime}$. By the above relation, both $H\circ\mathfrak{a}$ and $\mathfrak{a}^{\prime}\circ H$ lift to the same map and hence, $\mathfrak{a}^{\prime}=H\circ\mathfrak{a}\circ H^{-1}=\Phi_{H}^{+}(\mathfrak{a})$ and $A^{\prime}$ corresponds to a lift of $\Phi_{H}^{+}(\mathfrak{a})$.
A similar argument works for the lift of $\Phi_{H}^{-}(\mathfrak{a})$. ∎
Proof of Theorem 4.1 when $\frac{a}{d}$ is a $k$-th root of unity for some $k>0$
Note that
$\widetilde{H^{k}\circ\mathfrak{a}\circ H^{-k}}(z,\zeta)=\widetilde{{\Phi_{H}^{+}}^{k}(\mathfrak{a})}(z,\zeta)=(z+c_{0},\zeta)=A(z,\zeta).$
Hence $H^{k}\circ\mathfrak{a}=\mathfrak{a}\circ H^{k}$ and $H^{k}\circ\mathfrak{a}^{-1}=\mathfrak{a}^{-1}\circ H^{k}$ Now for $z\in K^{-}$, there exists $R_{z}\geq 0$ such that $\|H^{-n}(z)\|\leq R_{z}$. Since $\mathfrak{a}(\overline{B(0;R_{z})})$ is a compact subset of $\mathbb{C}^{2}$, there exists $M>0$ such that
$G_{H}^{-}(\mathfrak{a}(z))=d^{-n}G_{H}^{-}(\mathfrak{a}\circ H^{-n}(z))\leq d^{-n}M$
for large $n$. Thus $\mathfrak{a}(z)\in K^{-}$ and hence $\mathfrak{a}(K^{-})\subset K^{-}$. Similarly, by applying this to the automorphism $\mathfrak{a}^{-1}$, we get $\mathfrak{a}(K^{-})=K^{-}$. Hence by Theorem 1.1 in [5], $\mathfrak{a}=L\circ H^{s^{\prime}}$ where $L$ is an affine map of the form
$L(x,y)=(ex+f,e^{\prime}y+f^{\prime}),$
with $|e|=|e^{\prime}|=1.$ Also $G_{H}^{\pm}\circ L(x,y)=G_{H}^{\pm}(x,y)$ and $L(K^{\pm})=K^{\pm}$. For $(x,y)\in V^{+}\cap L^{-1}(V^{+})$
$|\phi(L(x,y))|=\exp({G_{H}^{+}(x,y)})=|\phi(x,y)|,$
where $\phi$ is the Böttcher function from Subsection 2.1. Since $V^{+}\cap L^{-1}(V^{+})$ is a non-empty open subset of $\mathbb{C}^{2}$ there exists $|\eta|=1$ such that
$\phi\circ L(x,y)=\eta\phi(x,y).$
for every $(x,y)\in V_{R}^{+}.$ In particular, $L^{*}\omega=\omega$ where $\omega=\frac{d\phi}{\phi}$. Thus $\alpha(L(C))=\alpha(C)$ for every curve $C\in U^{+}$ or $L\in{\rm Aut}_{1}(U^{+})$. Also, $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$ and hence from Proposition 2.3, $\mathfrak{a}=L$ and $s^{\prime}=0$.
Further note that $\mathfrak{a}\circ H^{k}=H^{k}\circ\mathfrak{a}$, and hence $e^{\prime d}=e^{\prime}$ and $e^{\prime d^{\prime}}=e$. Thus there are finitely many choices for $e^{\prime}$, and hence consequently, for $e$. For an integer $k\geq 1$, the locus of fixed points of $H^{k}$ forms a complex analytic subvariety that is compact since it is contained in both $K^{\pm}$, and hence $K$. By Bezout’s theorem, there are at most $d^{2k}$ fixed points of $H^{k}$. Also, $\mathfrak{a}$ should fix the set of fixed points of $H^{k}$. Hence there are finitely many choices of $f$ and $f^{\prime}$ for a pair $(e,e^{\prime})$ and this shows that there are only finitely many affine maps that can commute with $H^{k}$.
Remark 4.3.
The steps above also imply that if $f\in{\rm Aut}(\mathbb{C}^{2})\cap{\rm Aut}_{1}(U^{+})$ is such that $f\circ H^{m}=H^{m}\circ f$ for some $m\in\mathbb{Z}$, then $f(K^{\pm})=K^{\pm}$ and $f$ is an affine map. Conversely, if $f$ is an affine map preserving both $K^{\pm}$ then $f\in{\rm Aut}(\mathbb{C}^{2})\cap{\rm Aut}_{1}(U^{+})$.
Now recall that the $\mathfrak{G}$-lift of $\mathfrak{a}$ is $A(z,\zeta)=(z+c_{0},\zeta)$. Also, $\mathfrak{a}^{n}$ are automorphisms of $\mathbb{C}^{2}$ that $\mathfrak{G}$-lift to $A^{n}(z,\zeta)=(z+nc_{0},\zeta)$ for all $n\geq 1$. Since $\frac{d}{a}$ is a $k$-th root of unity by assumption, it follows that $H^{k}\circ\mathfrak{a}^{n}=\mathfrak{a}^{n}\circ H^{k}$. But there are only finitely many such maps that commute with $H^{k}$, and hence there exists $\tilde{n}>0$ such that $\mathfrak{a}^{\tilde{n}}=\textsf{Identity}.$ In particular, as $\mathfrak{G}$ is injective, the $\mathfrak{G}$-lift of $\mathfrak{a}^{\tilde{n}}$ is the identity. Hence $A^{\tilde{n}}(z,\zeta)=(z+\tilde{n}c_{0},\zeta)=(z,\zeta)$ or $c_{0}=0$. ∎
To complete the proof in the other cases, i.e., when $\left|\frac{d}{a}\right|=1$ but is not a root of unity and $\left|\frac{d}{a}\right|\neq 1$, several intermediate results will be needed.
Lemma 4.4.
Let $\{\mathfrak{a}_{n}\}\subset{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ be such that for every $n\geq 1$, the $\mathfrak{G}$-lift of $\mathfrak{a}_{n}$ is of the form
$A_{n}(z,\zeta)=(z+c_{n},\zeta)$
with $c_{n}\to c$ as $n\to\infty$. Then there exists $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ such that it $\mathfrak{G}$-lifts to $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ as
$A(z,\zeta)=(z+c,\zeta).$
Further, both $\mathfrak{a}_{n}\to\mathfrak{a}$, $\mathfrak{a}_{n}^{-1}\to\mathfrak{a}^{-1}$ on compact subsets of $\mathbb{C}^{2}$.
Proof.
Recall that $\widehat{\Pi}$ is the covering map from $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ to $U^{+}$ as in Theorem 2.1. Thus for every $z\in U^{+}$, there exists $\eta_{z}>0$ such that $\widehat{\Pi}$ restricted to each component of $\widehat{\Pi}^{-1}(B(z,\eta_{z}))$ is a biholomorphism. Also, from Step 4 in the proof of Theorem 3.4, there exists $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$ such that $\mathfrak{G}$-lift of $a$ is $A$.
Now fix a $z\in U^{+}$ and let $U$ be one such component of $\widehat{\Pi}^{-1}\big(B(\mathfrak{a}(z),\eta_{\mathfrak{a}(z)})\big)$ and $W=A^{-1}(U)$. There exists $\widetilde{W}\subset W$ such that $A_{n}(\widetilde{W})\subset U$ for every $n\geq n_{0}$ and
$\widetilde{W}\cap\widehat{\Pi}^{-1}(z)\neq\emptyset.$
Since $\widehat{\Pi}$ is continuous and the topologies on $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ and $U^{+}$ are equivalent to the topology induced by the standard norm of $\mathbb{C}^{2}$, it follows that for a given $\epsilon>0$ there exists $\delta>0$ such that
$|\widehat{\Pi}(w)-\widehat{\Pi}(\tilde{w}|<\epsilon\text{ whenever }|w-\tilde{w}|<\delta,$
for every $w,w^{\prime}\in U$. Furthermore, for $n\geq 1$ sufficiently large, note that $|A_{n}(w)-A(w)|<\delta$ for every $w\in\tilde{W}$, i.e., $|\widehat{\Pi}\circ A_{n}(w)-\widehat{\Pi}\circ A(w)|<\epsilon$. But $\widehat{\Pi}\circ A_{n}=\mathfrak{a}_{n}\circ\widehat{\Pi}$ and $\widehat{\Pi}\circ A=\mathfrak{a}\circ\widehat{\Pi}$. Thus
$|\mathfrak{a}_{n}\circ\widehat{\Pi}(w)-\mathfrak{a}\circ\widehat{\Pi}(w)|<\epsilon$
[…]
Similar arguments show that $\mathfrak{a}_{n}^{-1}\to A^{-1}$ on compact subsets of $U^{+}$. Next, we prove the following claim, which will complete the proof.
Claim: The maps $\mathfrak{a}^{\pm}$ extend as automorphisms of $\mathbb{C}^{2}$ and $\mathfrak{a}_{n}^{\pm}\to\mathfrak{a}^{\pm}$ on its compact subsets.
Let $(x_{0},y_{0})\in\mathbb{C}^{2}$. Consider the complex line $L_{x_{0}}=\{(x_{0},y):y\in\mathbb{C}\}$. Then $L_{x_{0}}\cap K^{+}$ is compact. Let $R_{x_{0}}=\max\{|x_{0}|+1,R_{H}\}>0$, where $R_{H}>0$ is the radius of filtration in (2.4) and note that $L_{x_{0}}\cap K^{+}$ is compactly contained in the disc
$\mathcal{D}_{x_{0}}=\{(x_{0},y):|y|<R_{x_{0}}\}.$
Let $D_{x_{0}}=L_{x_{0}}\setminus\mathcal{D}_{x_{0}}$ and for every $(x_{0},y)\in D_{x_{0}}$ define
(4.2) $\mathfrak{a}_{0}(x_{0},y)=\lim_{k\to\infty}\frac{1}{2\pi i}\int_{\gamma_{x_{0}}}\frac{\mathfrak{a}_{n}(x_{0},w)}{w-y}dw=\frac{1}{2\pi i}\int_{\gamma_{x_{0}}}\frac{\mathfrak{a}(x_{0},w)}{w-y}dw,$
where $\gamma_{x_{0}}=(x_{0},0)+(0,R_{x_{0}}e^{it})$, $t\in[0,2\pi]$. Since $\gamma_{x_{0}}$ is a compact subset of $U^{+}$ and the sequence of automorphisms $\mathfrak{a}_{n}\to\mathfrak{a}$, uniformly on compact subsets of $U^{+}$, $\mathfrak{a}_{0}(x_{0},y)$ is holomorphic in the $y$-variable on $D_{x_{0}}$. Also, $\mathfrak{a}_{0}(x_{0},y)=\mathfrak{a}(x_{0},y)$ on $\gamma_{x_{0}}$. Hence $\mathfrak{a}(x_{0},\cdot)$ extends as a holomorphic function on $L_{x_{0}}$. By repeating this reasoning for all $x$ sufficiently close to $x_{0}$ and noting that the same radius $R_{x_{0}}$ will work for such $x$, $\mathfrak{a}$ extends to a function on $\mathbb{C}^{2}$ that is holomorphic in the $y$-variable. Furthermore, by (4.2), $\mathfrak{a}_{n}$ converges uniformly on compact subsets of $V_{R_{H}}\cup V_{R_{H}}^{-}$ to $\mathfrak{a}_{0}$. In particular, the sequence of automorphisms $\{\mathfrak{a}_{n}\}$ converges uniformly on compact subsets of $\mathbb{C}^{2}$, and hence, $\mathfrak{a}$ extends to a holomorphic map on $\mathbb{C}^{2}$.
Finally, as $\mathfrak{a}$ is a limit of the automorphisms $\mathfrak{a}_{n}$ and $\mathfrak{a}$ restricted to $U^{+}$ is injective, $\mathfrak{a}$ is injective as well with $\mathfrak{a}(\mathbb{C}^{2})\subset\mathbb{C}^{2}$. Also, $\mathfrak{a}(U^{+})=U^{+}$ and $\mathfrak{a}(K^{+})\subset K^{+}$. Similar arguments (as above) applied to $\mathfrak{a}^{-1}$, show that $\mathfrak{a}^{-1}$ extends as an injective holomorphic map on $\mathbb{C}^{2}$ such that $\mathfrak{a}^{-1}(U^{+})=U^{+}$ and $\mathfrak{a}^{-1}(K^{+})\subset K^{+}$. Now $\mathfrak{a}\circ\mathfrak{a}^{-1}=\textsf{Identity}$ on $U^{+}$, hence on $\mathbb{C}^{2}$, i.e., $\mathfrak{a}^{\pm}(\mathbb{C}^{2})=\mathbb{C}^{2}$ and $\mathfrak{a}^{\pm}(K^{+})=K^{+}$. ∎
Lemma 4.5.
Suppose $\frac{d}{a}$ is not a root of unity. Further, let $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ be such that the $\mathfrak{G}$-lift of $\mathfrak{a}$ is of the form $A(z,\zeta)=(z+c_{0},\zeta)\text{ with }c_{0}\neq 0.$ Then,
  1. (i)
    $\mathbb{C}\leq{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ if $\frac{a}{d}\in\mathbb{C}\setminus\mathbb{R}$.
  2. (ii)
    $\mathbb{R}\leq{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ if $\frac{a}{d}\in\mathbb{R}$.
Proof.
For $n\geq 1$, define
[…]
Since the $\mathfrak{G}$-lifts of $\mathfrak{a}^{l}$, $l\in\mathbb{Z}$, are of the form
$A^{l}(z,\zeta)=(z+lc_{0},\zeta),$
the $\mathfrak{G}$-lifts of $\left(\Phi_{H}^{+}\right)^{n}\left(\mathfrak{a}^{l}\right)$ and $\left(\Phi_{H}^{-}\right)^{n}\left(\mathfrak{a}^{l}\right)$, $n\geq 1$, are of the form
[…]
respectively. For $n\in\mathbb{Z}$, define
$S_{n}=\{kc_{n}+lc_{n+1}:\text{ where }k,l\in\mathbb{Z}\}\text{ and }S=\bigcup_{n=-\infty}^{\infty}S_{n}.$
Note that the $\mathfrak{G}$-lift of $\left(\phi_{H}^{+}\right)^{n}(\mathfrak{a}^{k})\circ\left(\phi_{H}^{+}\right)^{n+1}(\mathfrak{a}^{l})\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ is of the form
$A_{n}^{k}\circ A_{n+1}^{l}(z,\zeta)=(z+kc_{n}+lc_{n+1},\zeta).$
Similarly, the $\mathfrak{G}$-lift of $\left(\phi_{H}^{-}\right)^{n}(\mathfrak{a}^{k})\circ\left(\phi_{H}^{-}\right)^{n-1}(\mathfrak{a}^{l})\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ is of the form
$\widetilde{A}_{n}^{k}\circ\widetilde{A}_{n-1}^{l}(z,\zeta)=(z+kc_{-n}+lc_{-n+1},\zeta).$
In particular, for every $\tilde{c}\in S$ there exists an element in $f\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ such that the $\mathfrak{G}$-lift of $f$ is $(z+\tilde{c},\zeta).$
Case 1: Let $\left|\frac{a}{d}\right|\neq 1$ and $\frac{a}{d}\in\mathbb{C}\setminus\mathbb{R}$.
Note that $S_{n}$ is a lattice of points generated by the complex numbers (or the vectors) $c_{n}$ and $c_{n+1}$ for every integer $n\in\mathbb{Z}$. The minimum distance between any two distinct points in this lattice is given by
$\min\{|p-q|:p,q\in S_{n}\text{ and }p\neq q\}\leq\left(\left|\frac{a}{d}\right|^{n+1}+\left|\frac{a}{d}\right|^{n}\right)|c_{0}|.$
If $\left|\frac{a}{d}\right|<1$, then for any point $p_{0}\in\mathbb{R}^{2}$
$\text{dist}(p_{0},S)\leq\text{dist}(p_{0},S_{n})<2\left|\frac{a}{d}\right|^{n}|c_{0}|\to 0\text{ as }n\to\infty.$
Otherwise, if $\left|\frac{a}{d}\right|>1$, by a similar argument as above, we see that
$\text{dist}(p_{0},S)\leq\text{dist}(p_{0},S_{-n})<2\left|\frac{d}{a}\right|^{n-1}|c_{0}|\to 0\text{ as }n\to\infty.$
Thus, $S$ is a dense subset of $\mathbb{R}^{2}\backsimeq\mathbb{C}$. Now, by Lemma 4.4, it follows that $\mathbb{C}\leq{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$.
Case 2: Let $\left|\frac{a}{d}\right|\neq 1$ and $\frac{d}{a}\in\mathbb{R}$. Then $S_{n}\subset\{tc_{0}:t\in\mathbb{R}\}$ and
$\min\{|p-q|:p,q\in S_{n}\text{ and }p\neq q\}\leq\left(\left|\frac{a}{d}\right|^{n+1}+\left|\frac{a}{d}\right|^{n}\right)|c_{0}|$
or
$\min\{|p-q|:p,q\in S_{-n}\text{ and }p\neq q\}\leq\left(\left|\frac{d}{a}\right|^{n-1}+\left|\frac{d}{a}\right|^{n}\right)|c_{0}|.$
Hence by the same reasoning as in Case 1, $S$ is a dense subset $\{tc_{0}:t\in\mathbb{R}\}.$ Thus by Lemma 4.4, it follows that $\mathbb{R}\leq{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$.
Case 3: Let $\left|\frac{a}{d}\right|=1$ and $\frac{a}{d}$ is not a root of unity.
Then for a positive integer $l\geq 1$, $\{lc_{n}\}_{n\in\mathbb{Z}}$ is a dense subset of the circle of radius $l|c_{0}|$. Hence, by Lemma 4.4, it follows that for every $\tilde{c}$ such that $|\tilde{c}|=l|c_{0}|$ there exists an element in $f_{\tilde{c}}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ such that the $\mathfrak{G}$-lift of $f_{\tilde{c}}$ is $(z+\tilde{c},\zeta).$ In particular, the $\mathfrak{G}$-lift of $\mathfrak{a}^{l}\circ f_{\tilde{c}}$ is
$(z+lc_{0}+\tilde{c},\zeta).$
Note that $0\leq|lc_{0}+\tilde{c}|\leq 2l|c_{0}|$, with the bounds being attained at the points $\tilde{c}=-lc_{0}$ and $\tilde{c}=lc_{0}$. Also, $lc_{0}+\tilde{c}$ is a continuous function on the circle of radius $l|c_{0}|$. Thus the $\mathfrak{G}$-lift of $\left(\Phi_{H}^{+}\right)^{n}(\mathfrak{a}^{l}\circ f_{\tilde{c}})\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ is of the form
$\Big(z+\left(\frac{a}{d}\right)^{n}(lc_{0}+\tilde{c}),\zeta\Big).$
[…]
Next, we complete the proof of Theorem 4.1 for the rest of the cases.
Proof of Theorem 4.1 when $\frac{d}{a}\in\mathbb{C}$ and not a root of unity
Suppose $c_{0}\neq 0$ in the assumption of Theorem 4.1. Then by Lemma 4.5, for every $r>0$ there exists $0<|\eta|<r$ and $f_{\eta}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ such that the $\mathfrak{G}$-lift of $f_{\eta}$ is $\tilde{f}_{\eta}(z,\zeta)=(z+\eta,\zeta)$ on $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$.
Recall the domains $\widetilde{W}_{M}^{+}\subset U^{+}$ and $\widetilde{\Omega}_{M}^{+}\subset\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ from the Subsection 2.2 and the biholomorphism $\widetilde{\Psi}$ between them. See also Figure 1 in this context. Note that for every $\epsilon>0$ there exists $\widetilde{R}_{\epsilon}^{1}>0$ such that for $|y|>2\widetilde{R}_{\epsilon}^{1}>0$, sufficiently large
$\widetilde{\Psi}(0,y)\in D(0;\epsilon)\times D(0;\widetilde{R}_{\epsilon}^{1})^{c}.$
By Lemma 2.9 for $\epsilon>0$ there exists $\widetilde{R}_{\epsilon}^{2}>0$ such that
$D(0;2\epsilon)\times D(0;\widetilde{R}_{\epsilon}^{2})^{c}\subset\widetilde{\Omega}_{M}^{+}.$
Thus for a given $\epsilon>0$ there exists $\widetilde{R}_{\epsilon}>\max\{\widetilde{R}_{\epsilon}^{1},\widetilde{R}_{\epsilon}^{2}\}+\epsilon$ such that
$\left[\widetilde{\Psi}\left(\{0\}\times D\left(0;\widetilde{R}_{\epsilon}\right)^{c}\right)\right]_{\epsilon}\subset\widetilde{\Omega}_{M}^{+},$
where $X]_{\epsilon}$ denote the $\epsilon$-neighborhood of a set $X$. Now, again choose $0<|\eta|<\epsilon$ such that $f_{\eta}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ and $\mathfrak{G}$-lift of $f_{\eta}$ is $\tilde{f}_{\eta}(z,\zeta)=(z+\eta,\zeta)$. Thus for $|y|>2\widetilde{R}_{\epsilon}^{c}$,
(4.3) $f_{\eta}(0,y)=\widetilde{\Psi}^{-1}\circ\tilde{f}_{\eta}\circ\widetilde{\Psi}(0,y)\in D(0;2\epsilon)\times D\left(0;\widetilde{R}_{\epsilon}\right)^{c}\subset\widetilde{W}_{M}^{+}.$
Let $f_{\eta}(0,y)=(p_{\eta}(y),q_{\eta}(y))$. Since $f_{\eta}\in{\rm Aut}(\mathbb{C}^{2})$, both $p_{\eta},q_{\eta}$ are entire functions on $\mathbb{C}$. Hence from (4.3), we have
(4.4) $\widetilde{\Psi}\circ(p_{\eta}(y),q_{\eta}(y))=\tilde{f}_{\eta}\circ\widetilde{\Psi}(0,y).$
Let us also recall the map $\widetilde{\Psi}$ explicitly from subsection 2.2:
(4.5) $\widetilde{\Psi}(x,y)=\Big(\psi\big(x,\phi(x,y)\big)-R\big(\phi(x,y)\big),\phi(x,y)\Big),\text{ where }\psi(x,y)=y\int_{0}^{x}\frac{\partial\lambda}{\partial y}(t,y)dt,$
where $\phi$ is the Böttcher coordinate function defined in (2.5) and $\lambda$, $R$ are the functions discussed in Lemma 2.4, Remark 2.5 and Proposition 2.8 respectively. Then from (4.4), it follows that
$\phi(p_{\eta}(y),q_{\eta}(y))=\phi(0,y).$
By equating the second coordinate,
$\psi(0,\phi(0,y))-R(\phi(0,y))+\eta$ $=\psi\big(p_{\eta}(y),\phi(p_{\eta}(y),q_{\eta}(y))\big)-R\left(\phi\big(p_{\eta}(y),q_{\eta}(y)\big)\right)$
$=\psi\big(p_{\eta}(y),\phi(0,y)\big)-R(\phi(0,y)).$
Also note that $\psi(0,y)=0$ and hence the above reduces to
$\eta=\psi\big(p_{\eta}(y),\phi(0,y)\big)=\phi(0,y)\int_{0}^{p_{\eta}(y)}\frac{\partial\lambda}{\partial y}\big(t,\phi(0,y)\big)dt.$
By the estimates on $\phi$ and $\frac{\partial\lambda}{\partial y}(t,y)$ obtained in Lemma 2.4 and Remark 2.5
$\eta-yp_{\eta}(y)$ $=\phi(0,y)\int_{0}^{p_{\eta}(y)}\frac{\partial\lambda}{\partial y}\big(t,\phi(0,y)\big)dt-y\int_{0}^{p_{\eta}(y)}dt$
$=\left(\frac{\phi(0,y)}{y}-1\right)y\int_{0}^{p_{\eta}(y)}\frac{\partial\lambda}{\partial y}\big(t,\phi(0,y)\big)dt+y\int_{0}^{p_{\eta}(y)}\left(\frac{\partial\lambda}{\partial y}\big(t,\phi(0,y)\big)-1\right)dt.$
For given $\delta>0$ sufficiently small there exists $\widetilde{R}_{\delta}>\widetilde{R}$, sufficiently large, such that $|y|>\widetilde{R}_{\delta}$
$\left|\frac{\eta}{y}-p_{\eta}(y)\right|\leq\delta(1+\delta)|p_{\eta}(y)|+\delta|p_{\eta}(y)|=(\delta^{2}+2\delta)|p_{\eta}(y)|.$
Let $\delta=\frac{1}{4}$. Then for $|y|>\widetilde{R}_{1/4}$, $|p_{\eta}(y)|\leq\frac{16}{7}\frac{|\eta|}{|y|}.$ In particular, $p_{\eta}$ is an entire function and $p_{\eta}(y)\to 0$ as $y\to\infty$. Hence $p_{\eta}(y)\equiv 0$. Thus from the above, $\eta=\psi(0,\phi(0,y))=0$, which is a contradiction to the assumption. This establishes that $c_{0}=0$ and completes the proof of Theorem 4.1. ∎
Remark 4.6.
It should be noted that the implications of Lemma 4.5 as well as Lemma 4.4 could have been avoided to complete the proof of Theorem 4.1. However, understanding them provides additional information and helps to establish the contradiction.
4.2. The group ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$
The goal here is to further develop the observation in Theorem 4.1 to prove the following result.
Theorem 4.7.
The group ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ is a subgroup of $\mathbb{Z}_{d_{0}(d-1)}$ where $d_{0}$ is as in (3.5). In particular, it is the set of all the affine maps that preserve both $K^{\pm}$.
Proof.
Suppose not. Let the number of elements in ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ be more than $d_{0}(d-1)$. Recall the map $\mathfrak{G}$ from Step 5 in the proof of Theorem 3.1. Then there exist distinct $f,g\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ such that
$\mathfrak{G}(f)=(c_{1},\alpha)\text{ and }\mathfrak{G}(g)=(c_{2},\alpha)$
where $\alpha$ is a $d_{0}(d-1)$-root of unity. Thus, $\mathfrak{G}(f\circ g^{-1})=(\tilde{c},1)$, or equivalently the $\mathfrak{G}$-lift of $f\circ g^{-1}$ is of the form $(z+\tilde{c},\zeta)$. Since $f\circ g^{-1}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$, by Theorem 4.1, $\tilde{c}=0$, i.e., $f\circ g^{-1}=\textsf{Identity}$ or $f=g$. This is a contradiction.
Now suppose $f\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ then $\left(\Phi_{H}^{+}\right)^{n}(f)\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ for every $n\geq 1$. Since there are only finitely many maps in ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$, there exist distinct $n,m\geq 1$ such that
$\left(\Phi_{H}^{+}\right)^{n}(f)=\left(\Phi_{H}^{+}\right)^{m}(f)\text{ or }H^{n-m}\circ f=f\circ H^{n-m}.$
Thus, from Remark 4.3, $f$ is an affine map that preserves both $K^{\pm}$. Also note $f^{n}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ for every $n\geq 1$. Thus, there exists $1\leq n_{f}\leq d_{0}(d-1)$ such that $f^{n_{f}}=\textsf{Identity}.$
Now let $L$ be an affine map preserving $K^{+}$. Then for every $n\geq 1$
$L\circ H^{n}\circ L^{-1}\circ H^{-n}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2}).$
As before there exist distinct integers $m,n\geq 1$ such that
$L\circ H^{n}\circ L^{-1}\circ H^{-n}=L\circ H^{m}\circ L^{-1}\circ H^{-m}\text{ or }H^{m-n}\circ L=L\circ H^{m-n}.$
Hence $L$ preserves both $K^{\pm}$ and the proof follows from the converse part of Remark 4.3. ∎
Now by exactly repeating the arguments in the later part of the above proof completes the
Proof of Theorem 1.1.
Let $f\in{\rm Aut}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2}).$ Then, as before, for every $n\geq 1$
$f\circ H^{n}\circ f^{-1}\circ H^{-n}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2}).$
In particular, there exist distinct integers $m,n\geq 1$ such that
$f\circ H^{n}\circ f^{-1}\circ H^{-n}=f\circ H^{m}\circ f^{-1}\circ H^{-m}\text{ or }H^{m-n}\circ f=f\circ H^{m-n}.$
Thus, $f$ preserves both $K^{\pm}$. Hence by Theorem 1.1 in [5], $f=L\circ H^{s^{\prime}}$ where $L$ is an affine map that preserves both $K^{\pm}$. ∎
We conclude this section with two immediate corollaries. First, the very rigid property of the set $K^{+}$ allows us to weaken the assumptions of Theorem 1.1, stated as
Corollary 4.8.
Let $H$ be a generalised Hénon map of the form (1.1) and $\phi\in{\rm Aut}(\mathbb{C}^{2})$ such that $\phi(K_{H}^{+})\subset K_{H}^{+}$. Then $\phi=L\circ H^{s}$ where $L$ is a linear map preserving both $K_{H}^{\pm}$ and $s$ is an integer.
Proof.
Recall that $\mu_{H}^{+}=\frac{1}{2\pi}dd^{c}G_{H}^{+}$. Note that $\phi_{*}(\mu_{H}^{+})$ is positive $dd^{c}$-closed $(1,1)$ current supported in $K_{H}^{+}$, hence by the uniqueness of closed positive currents supported in $K_{H}^{\pm}$ — see [13], [10] — $\phi_{*}(\mu_{H}^{+})=c\mu_{H}^{+}$ for some $c>0$. Thus, by the proof Theorem A of [9], $G_{H}^{+}\circ\phi=cG_{H}^{+}$. Hence $\phi(K_{H}^{+})=K_{H}^{+}$, and the result follows from Theorem 1.1. ∎
Finally, we use the above to prove a version of [9, Theorem A] under comparatively weaker assumptions.
Corollary 4.9.
Let $H$ and $F$ be generalised Hénon maps of the form (1.1) and assume that there exists an automorphism $\phi$ satisfying $\phi(K_{F}^{+})\subset K_{H}^{+}$. Then $\phi$ is a polynomial automorphism and $H$ and $F$ are polynomially conjugated in $\mathbb{C}^{2}$.
Proof.
Note that, $\phi^{-1}\circ H\circ\phi(K_{F}^{+})\subset K_{F}^{+}$ and is also an automorphism of $\mathbb{C}^{2}$. So by the above result, there exist an affine map $L$ and $s\in\mathbb{Z}$ such that $\phi^{-1}\circ H\circ\phi=L\circ F^{s}$. In particular,
$L^{-1}\circ\phi^{-1}\circ H\circ\phi\circ L=F^{s}\circ L,\text{ i.e., }(\phi\circ L)^{-1}\circ H\circ(\phi\circ L)=F^{s}\circ L.$
Also, $L(K_{F}^{\pm})=K_{F}^{\pm}$ and there exists $k\geq 1$ such that $F^{k}\circ L=L\circ F^{k}$. Then there exist a subsequence $\{n_{k}\}$ of positive integers and a positive integer $n_{0}$ such that
$(\phi\circ L)^{-1}\circ H^{n_{k}}\circ(\phi\circ L)=F^{sn_{k}}\circ L^{n_{0}}.$
Now for $z\in K_{F}^{+}\setminus K_{F}$, the left hand side of the above identity is bounded, however the right side diverges if $s$ is a negative non-zero integer. Hence $s\geq 0$. By Theorem A of [9], it follows that $\phi\circ L$ is a polynomial automorphism, and consequently $\phi$ is a polynomial. ∎
5. Proof of Theorem 1.2 and Corollary 1.3
Recall that for $c>0$,
$\Omega_{c}=\{z\in\mathbb{C}^{2}:G_{H}^{+}(z)<c\}\text{ and }\Omega_{c}^{\prime}=\Omega_{c}\setminus K^{+}$
Consider the group isomorphisms $\Phi_{H,c}^{\pm}:{\rm Aut}(\Omega_{c}^{\prime})\to{\rm Aut}(\Omega_{cd^{\pm}}^{\prime})$ on ${\rm Aut}(\Omega_{c}^{\prime})$, analogous to the definitions in (4.1).
(5.1) $\Phi_{H,c}^{+}(f_{c})$ $=H\circ f_{c}\circ H^{-1}\text{ and }\Phi_{H,c}^{-}(f_{c})=H^{-1}\circ f_{c}\circ H\text{ for }f_{c}\in{\rm Aut}(\Omega_{c}^{\prime}).$
for every $c>0$. As before,
$\Phi_{H,c}^{\pm}\left({\rm Aut}_{1}(\Omega_{c}^{\prime})\right)={\rm Aut}_{1}(\Omega_{cd^{\pm}}^{\prime}).$
Proof of Theorem 1.2.
For $\mathfrak{a}\in{\rm Aut}_{1}(U^{+})$, by Theorem 3.1, there exists a unique $\mathfrak{G}$-lift of the form $\tilde{\mathfrak{a}}=(\beta z+\gamma(\zeta),\alpha\zeta)$, such that $\alpha^{d_{0}(d-1)}=1$ and $\beta=\alpha^{d+d^{\prime}}$. Note $\tilde{\mathfrak{a}}(\mathbb{C}\times\mathcal{A}_{c})=\mathbb{C}\times\mathcal{A}_{c}$ for every $c>0$. Hence, by Theorem 3.2, $\mathfrak{a}(\Omega_{c}^{\prime})=\Omega_{c}^{\prime}$ for every $c>0$ and $\mathfrak{a}$ induces identity in the fundamental group of $\Omega_{c}^{\prime}$, i.e., $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ and ${\rm Aut}_{1}(U^{+})\subset{\rm Aut}_{1}(\Omega_{c}^{\prime})$ for every $c>0$.
Recall the map $\mathfrak{G}_{c}$ discussed in the proof Theorem 3.2, depending on the unique lift of form (3.7) of every element in ${\rm Aut}_{1}(\Omega_{c}^{\prime})$. As in earlier sections, we will denote this lift as the $\mathfrak{G}_{c}$-lift. By Theorem 3.2, the covering map $\widehat{\Pi}$ is the same on every $\mathbb{C}\times\mathcal{A}_{c}$, where $c>0$. Suppose $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ is such that $\mathfrak{G}_{c}$-lift of $\mathfrak{a}$ is $(z+\eta,\zeta)$. Now by Remark 3.3, $(z+\eta,\zeta)$ also corresponds to the $\mathfrak{G}$-lift of an element $\mathfrak{a}^{\prime}\in{\rm Aut}_{1}(U^{+})$. Thus, by Theorem 3.2, on $\mathbb{C}\times\mathcal{A}_{c}$
[…]
on $\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$. In particular, $\mathfrak{a}^{\prime}$ is the extension of the map $\mathfrak{a}$ from $\Omega_{c}^{\prime}$ to $U^{+}$.
Suppose $\mathfrak{a}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$ is such that $\mathfrak{G}_{c}$-lift of $\mathfrak{a}$ is
$(\alpha^{d+d^{\prime}}z+\gamma(\zeta),\alpha\zeta)$
where $\alpha\in\mathbb{Z}_{d_{0}(d-1)}$ and $\alpha\neq 1$. Let $\mathfrak{a}_{n}=(\Phi_{H,c}^{+})^{n}\in{\rm Aut}_{1}(\Omega^{\prime}_{cd^{n}})$ for every $n\geq 1$. Furthermore, let the $\mathfrak{G}_{cd^{n}}$-lift of $\mathfrak{a}_{n}$ be of the form $(a_{n},\alpha_{n})$. Then for some positive integers $m>n\geq 1$, $\alpha_{n}=\alpha_{m}$. In particular as $m>n$, $\mathfrak{a}_{n}\circ\mathfrak{a}_{m}^{-1}\in{\rm Aut}_{1}(\Omega^{\prime}_{cd^{n}})$ and the $\mathfrak{G}_{cd^{n}}$-lift of $\mathfrak{a}_{n}\circ\mathfrak{a}_{m}^{-1}$ is of the form $(z+\eta_{nm},\zeta)$, $\eta_{nm}\in\mathbb{C}$, by similar steps as in Step 4 of Theorem 3.1 and Remark 3.3. Now, by the above, $f=\mathfrak{a}_{n}\circ\mathfrak{a}_{m}^{-1}$ extends to $U^{+}$ and is in ${\rm Aut}_{1}(U^{+})$. Consequently, $f\in{\rm Aut}_{1}(\Omega_{cd^{m}}^{\prime})$ by the first part of this proof. Hence,
$\mathfrak{a}_{n}=f\circ\mathfrak{a}_{m}\in{\rm Aut}_{1}(\Omega_{cd^{m}}^{\prime}).$
Thus $\mathfrak{a}_{n}$ extends from $\Omega_{cd^{n}}^{\prime}$ to $\Omega_{cd^{m}}^{\prime}$, or equivalently $\mathfrak{a}=(\Phi_{H}^{-})(\mathfrak{a}_{n})$ extends from $\Omega_{c}^{\prime}$ to $\Omega_{cd^{m-n}}^{\prime}$.
Repeating the above argument inductively, we get that $\mathfrak{a}$ extends to an automorphism of $\Omega_{cd^{k(m-n)}}^{\prime}$ for every $k\geq 1$. Now $U^{+}=\cup_{k=1}^{\infty}\Omega_{cd^{k(m-n)}}^{\prime}$. Thus, $\mathfrak{a}$ extends to an element of ${\rm Aut}_{1}(U^{+})$ or ${\rm Aut}_{1}(\Omega_{c}^{\prime})\subset{\rm Aut}_{1}(U^{+}).$
Next as an application of Theorem 1.2, we will complete
Proof of Corollary 1.3.
Note that by [6, Proposition 1.2], $f(K^{+})=K^{+}$, $f(\Omega_{b})=\Omega_{b}$ for $0<b<c$ and $f\in{\rm Aut}(\Omega_{c}^{\prime})$. For every $n\geq 1$ let
$\tilde{f}_{n}=f\circ H^{n}\circ f^{-1}\circ H^{-n}\text{ on }\Omega_{c}.$
Then $\tilde{f}_{n}\in{\rm Aut}_{1}(\Omega_{c}^{\prime})$. By Theorem 1.2, $\tilde{f}_{n}$ extends to an automorphism of $U^{+}$ that induces identity on the fundamental group of $U^{+}$. Also, as $\tilde{f}_{n}$ is holomorphic on a neighbourhood of $K^{+}$, $\tilde{f}_{n}$ extends as an automorphism of $\mathbb{C}^{2}$, say $F_{n}$. Thus, by Theorem 4.1, $F_{n}=L_{n}$ where $L_{n}$ is an affine map preserving $K^{+}$. By Theorem 1.1, there are only finitely many choices for every $L_{n}$, $n\geq 1$. Also, $f$ extends to an automorphism of $\Omega_{cd^{n}}$ as
$f^{-1}=H^{n}\circ f^{-1}\circ H^{-n}\circ L_{n}^{-1}\text{ on }\Omega_{cd^{n}}.$
Thus $f\in{\rm Aut}(\mathbb{C}^{2})$ as $\mathbb{C}^{2}=\cup_{n}\Omega_{cd^{n}}$, with $f(K^{+})=K^{+}$. Thus by Theorem 1.1, $f=L\circ H^{s}$ where $s\in\mathbb{Z}$ where $L$ is an affine map preserving $K^{\pm}$. Further $f(\Omega_{c})=\Omega_{c}$, hence $f=L$. ∎
Finally, we conclude with the following application on biholomorphic Short $\mathbb{C}^{2}$’s
Theorem 5.1.
Let $H$ and $F$ be generalised Hénon maps of the form (1.1) of the same degree $d\geq 2$ and assume that $d=\mathbf{p}^{m}$ for some prime $\mathbf{p}$ and $m\geq 1$. Suppose that $\Omega_{H,c}$ and $\Omega_{F,c^{\prime}}$ are biholomorphic, where
$\Omega_{H,c}=\{z\in\mathbb{C}^{2}:G_{H}^{+}(z)<c\}\text{ and }\Omega_{F,c^{\prime}}=\{z\in\mathbb{C}^{2}:G_{F}^{+}(z)<c^{\prime}\},$
for some $c,c^{\prime}>0$. Then there exist affine maps $A_{1}$ and $A_{2}$ such that
$H=A_{1}\circ F\circ A_{2}.$
Proof.
Let $K_{H}^{+}$ and $K_{F}^{+}$ denote the non-escaping set of $F$ and $H$, respectively. Let $p$ be a saddle point of $H$ such that the stable manifold $W^{s}(p)$ is dense in $J_{H}^{+}$. Since $J_{H}^{+}=\partial K_{H}^{+}$, $J_{F}^{+}=\partial K_{F}^{+}$ are non-smooth (see [2]), $G_{F}^{+}\circ\phi(W^{s}_{p})$ is bounded and subharmonic, i.e., it is constant. Hence, $\phi(J_{H}^{+})\subset K_{F}^{+}$. Similarly, $\phi^{-1}(J_{F}^{+})\subset K_{H}^{+}$. Now by following the arguments as in the proof of Proposition 1.2 in [6], $G_{F}^{+}\circ\phi(\partial\Omega_{b})=a_{b}>0$ for every $0<b<c$. Hence the punctured short $\mathbb{C}^{2}$’s are biholomorphic, i.e.,
$\Omega_{H,c}\setminus K_{H}^{+}=\Omega_{H,c}^{\prime}\simeq\Omega_{F,{c^{\prime}}}^{\prime}=\Omega_{F,{c^{\prime}}}\setminus K_{F}^{+}.$
The fundamental groups of the punctured Short $\mathbb{C}^{2}$’s are isomorphic, in particular,
$\mathbb{Z}\left[\frac{1}{d}\right]\simeq\uppi_{1}\left(\Omega_{H,c}^{\prime}\right)\simeq\uppi_{1}\left(\Omega_{F,{c^{\prime}}}^{\prime}\right).$
Let $\phi:\Omega_{F,c^{\prime}}\to\Omega_{H,c}$ be the biholomorphism. Define
(5.2) $\mathfrak{a}_{n}=\phi^{-1}\circ F^{n}\circ\phi\circ H^{-n},$
for every $n\geq 1$. Note that each $\mathfrak{a}_{n}$ induces the identity map on the fundamental group of $\Omega_{H,cd^{n}}$, i.e., $\mathfrak{a}_{n}\in{\rm Aut}_{1}(\Omega_{H,cd^{n}}^{\prime})$ for every $n\geq 1$. By Theorem 1.2, every $\mathfrak{a}_{n}$ extends to an automorphism of $U_{H}^{+}$, that induces identity on the fundamental group of $U_{H}^{+}$. Hence $\mathfrak{a}_{n}\in{\rm Aut}_{1}(U_{H}^{+})\cap{\rm Aut}(\mathbb{C}^{2})$ and $\mathfrak{a}_{n}(K_{H}^{+})=K_{H}^{+}$. Further, the above identity (5.2), now can be used to extend $\phi$ from $\Omega_{F,d^{n}c^{\prime}}$ to $\Omega_{H,cd^{n}}$, via the relation
$\phi=F^{n}\circ\phi\circ H^{-n}\circ\mathfrak{a}_{n}^{-1}.$
Thus $\phi$ extends to an automorphism of $\mathbb{C}^{2}$ such that $\phi(K_{F}^{+})=K_{H}^{+}$, or equivalently their corresponding escaping sets $U_{F}^{+}$ and $U_{H}^{+}$ are biholomorphic, via the map $\phi$. In particular by Theorem 1.1, each $\mathfrak{a}_{n}$, $n\geq 1$ is an affine map such that $\mathfrak{a}_{n}(K_{H}^{-})=K_{H}^{-}$, $G_{H}^{-}\circ{\mathfrak{a}}_{n}=G_{H}^{-}$ and $\mathfrak{a}_{n}\in\mathcal{L}$. Since there are only finitely many choice for every $\mathfrak{a}_{n}$, it follows from (5.2) that $\phi(K_{H}^{-})=K_{F}^{-}$ or $\phi(U_{H}^{-})=U_{F}^{-}$.
Thus $\phi^{*}(\mu_{H}^{-})$ and $\phi^{*}(\mu_{H}^{+})$ are $dd^{c}$-closed $(1,1)$-positive currents supported on $K_{H}^{-}$ and $K_{H}^{-}$, respectively. By arguing similarly as in the proof of Theorem A in [9, page 3748], there exist constants $a,b>0$ such that
$G_{F}^{+}\circ\phi(z)=aG_{H}^{+}(z)\text{ and }G_{F}^{-}\circ\phi(z)=bG_{H}^{-}(z).$
Since $\phi$ is an automorphism of $\mathbb{C}^{2}$,
$1=\int_{\mathbb{C}^{2}}\mu_{H}^{+}\wedge\mu_{H}^{-}=\int_{\phi(\mathbb{C}^{2})}\phi^{*}(\mu_{H}^{+}\wedge\mu_{H}^{-})=\int_{\mathbb{C}^{2}}ab\mu_{F}^{+}\wedge\mu_{F}^{-}=ab.$
Hence $b=a^{-1}$.
Claim: There exists $s\in\mathbb{Z}$ such that
[…]
$f\circ\widehat{\Pi}_{H}=\widehat{\Pi}_{F}\circ\tilde{f}$.
Again by the arguments as in Step 1 of the proof of Theorem 3.1,
$\tilde{f}(z,\zeta)=(\beta(\zeta)(z)+\gamma(\zeta),\alpha\zeta)$
where $|\alpha|=1$ and $\beta$, $\gamma$ holomorphic function on $\mathbb{C}\setminus\overline{\mathbb{D}}$. Thus if $(x,y)\in\Omega_{c,H}$. Let $(z,\zeta)\in\mathbb{C}\times\left(\mathbb{C}\setminus\overline{\mathbb{D}}\right)$ such that $\widehat{\Pi}_{H}(z,\zeta)=(x,y)$. Then $\widehat{\Pi}_{F}\circ\tilde{f}(z,\zeta)\in\Omega_{c,F}^{\prime}$, or $f(\Omega_{c,H}^{\prime})\subset\Omega_{c,F}^{\prime}$. Similarly for $f^{-1}$, hence $f(\Omega_{c,H}^{\prime})=\Omega_{c,F}^{\prime}$. Thus the claim.
Thus $G_{F}\circ f=G_{H}$ where $G_{H}=\max\{G_{H}^{+},G_{H}^{-}\}$ and $G_{F}=\max\{G_{F}^{+},G_{F}^{-}\}$. Now for $p\in\mathbb{C}^{2}$
$G_{F}(p)=\log|p|+O(1)\text{ and }G_{H}(f(p))=\log|f(p)|+O(1)=\log|p|+O(1).$
Hence $\|f(p)\|<M\|p\|+C$ for every $p\in\mathbb{C}$. This proves that $f$ is an affine map. Further, from (5.2), if $s\leq 0$,
$\mathfrak{a}_{-s+1}=\phi^{-1}\circ F^{-s+1}\circ\phi\circ H^{s-1}\text{ and }F^{s}\circ\phi=\phi\circ H^{s}\circ\mathfrak{a}_{-s}^{-1}.$
Thus $F^{s}\circ\phi$ is an affine map. Finally,
$H=\left(F^{s}\circ\phi\right)^{-1}\circ F\circ\left(\phi\circ H^{s}\circ\mathfrak{a}_{-s+1}^{-1}\right),$
which completes the proof. Since $\phi$ extends to an automorphism, (5.2) is true for $n\leq 0$. Hence, a similar argument works if $s>0$. ∎
Remark 5.2.
Note that the isomorphism induced by the map $\phi$ (in the above proof) on the fundamental group $\mathbb{Z}\left[\frac{1}{d}\right]$, is $\mathfrak{i}_{\phi}(x)=u_{\phi}x$, where $u_{\phi}$ is an unit in the ring $\mathbb{Z}\left[\frac{1}{d}\right]$. Suppose $u_{\phi}d^{n}\neq 1$ for every $n\in\mathbb{Z}$, then the above argument will not work, hence Theorem 5.1 could not be stated generally for arbitrary degrees $d\geq 2$.